### v1

# A Pre-trained Sequential Recommendation Framework: Popularity Dynamics for Zero-shot Transfer

Abstract.
Sequential recommenders are crucial to the success of online applications, e.g., e-commerce, video streaming, and social media. While model architectures continue to improve, for every new application domain, we still have to train a new model from scratch for high quality recommendations. On the other hand, pre-trained language and vision models have shown great success in zero-shot or few-shot adaptation to new application domains. Inspired by the success of pre-trained models in peer AI fields, we propose a novel pre-trained sequential recommendation framework: PrepRec. We learn universal item representations by modeling item popularity dynamics. Through extensive experiments on five real-world datasets, we show that PrepRec, without any auxiliary information, can not only zero-shot transfer to a new domain, but achieve competitive performance compared to state-of-the-art sequential recommender models with only a fraction of the model size. In addition, with a simple post-hoc interpolation, PrepRec can improve the performance of existing sequential recommenders on average by 13.8% in Recall@10 and 29.5% in NDCG@10. We provide an anonymized implementation of PrepRec at https://anonymous.4open.science/r/PrepRec--2F60/.
Recommender System, Neural Collaborative Filtering, Sequential Recommendation, Zero-shot Sequential Recommendation
copyright: acmcopyrightjournalyear: 2018doi: 10.1145/1122445.1122456copyright: noneconference: Woodstock ’18: ACM Symposium on Neural Gaze Detection; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-$9$-9/18/06
1. Introduction
Modeling sequential user behavior is critical to the success of online applications such as e-commerce, video streaming, and social media. Despite essential innovations for tackling the sequential recommendation task (Rendle et al., 2010; Hidasi et al., 2015; Tan et al., 2016; Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020), these models have to be trained from scratch for each application domain. because they learn representations for each item (Sun et al., 2019; Kang and McAuley, 2018) in the domain. Further, to maintain high performance, we need to retrain the models when there is a large influx of new users and items. Since the models learn representations for each item, they cannot generalize to new domains. Thus, in this paper, we ask: can we build a pre-trained sequential recommender system capable of cross-domain and cross-application zero-shot transfer without any auxiliary information? (e.g., using the model trained for online shopping in the US to predict the next movie that a user in India will watch).
At first glance, developing pre-trained sequential recommenders for zero-shot cross-domain inference seems impossible. While we see pre-trained language (Devlin et al., 2018b; Brown et al., 2020; OpenAI, 2023; Raffel et al., 2020; Liu et al., 2019) and vision models (Dosovitskiy et al., 2020; He et al., 2016; Radford et al., 2021) show excellent generalizability across datasets and applications, being able to achieve state-of-the-art performance by just a few fine-tuning steps (Devlin et al., 2018b; Liu et al., 2019) or even without any training (OpenAI, 2023; Brown et al., 2020) (i.e., zero-shot transfer), there are are essential differences. The representations learned by the pre-trained language model seem universal since the training domain and the application domain (e.g., text prediction and generation) share the same language and vocabulary, supporting the effective reuse of the word representations. However, in the cross-domain recommendation problem, the items are distinct across domains in sequential recommendation datasets (e.g., grocery items vs movies). Therefore, forming such generalizable correspondence is nearly impossible if we learn specific representations for each item within each domain. Recent work explores pre-trained models for sequential recommendation (Ding et al., 2021; Hou et al., 2022; Hao et al., 2021) within the same application (e.g., online retail). However, they assume access to auxiliary information of items (e.g., item description), which is application-dependent and is not always available in practice. Therefore, these models cannot learn universal representations of items; instead, they bypass the representation learning problem by using additional item side information. Thus, while achieving success within a domain, these recent models cannot generalize to new domains without auxiliary information. This paper tackles the challenge of zero-shot, cross-domain sequential recommendations without any auxiliary information.
Present Work: Our crucial insight is to learn the universal item representations based on the item popularity dynamics and use these representations for cross-domain, zero-shot prediction. We know that the marginal distribution of user and item activities is universal across datasets, supported by prior work in network science (Barabási and Albert, 1999; Barabasi, 2005) and by experiments in recommender systems (Salganik et al., 2006). In addition, recent work in recommender systems suggests that the popularity dynamics of items are also crucial for predicting users’ behaviors (Ji et al., 2020). Therefore, in this paper, we explicitly model the popularity dynamics of items and propose a novel pre-trained sequential recommendation framework: PrepRec. We learn to represent items based on their popularity dynamics, instead of their explicit item IDs. We encode the time interval between two consecutive interactions via relative-time encoding and the position of each interaction in the sequence by positional encoding. In contrast, prior work in sequential recommendation uses logical indices. Using physical time ensures that the predictions are not anti-causal, i.e., using the future interactions to predict the present. We learn the representations within a transformer architecture coupled with optimizations. The learned item popularity representations, time-interval and positional encoding are all universal, making it possible to build a pre-trained sequential recommender system capable of cross-domain and cross-application transfer without any auxiliary information. Our key contributions are as follows:
Universal item representations::
We are the first to learn universal item representations for sequential recommendation. In contrast, prior research learns item representations for each item ID or through item auxiliary information. We learn universal item representations by exploiting item popularity dynamics. We learn two temporal representations using a transformer architecture with optimizations at any time $t$ for each item’s popularity: at a coarse and fine-grained level. We represent items’ popularity dynamics (i.e., representing popularity changes) by concatenating representations over a fixed time interval. Item dynamics are inferrable from the user-item interaction data, and thus, the learned item representations are transferable across domains and applications. These item representations make possible pre-trained sequential recommender systems capable of cross-domain and cross-application transfer without any auxiliary information.
Zero-shot transfer without auxiliary information::
We propose a new challenging setting for pre-trained sequential recommender systems: zero-shot transfer without any auxiliary information. In contrast, previous works in sequential recommender systems capable of cross-domain zero-shot rely heavily on application-dependent auxiliary information (Ding et al., 2021; Hou et al., 2022; Hao et al., 2021). To the best of our knowledge, we are the first to tackle this challenging setting in sequential recommendation.
With extensive experiments, we empirically show that PrepRec has excellent generalizability across domains and applications. Remarkably, had we trained a state of the model from scratch for the target domain, instead of zero-shot transfer using PrepRec , the performance gain would have been only (on average) 0.2%. In addition, we show that with a post-hoc interpolation, PrepRec can outperform the state-of-the-art sequential recommender system on average by 13.8% in Recall@10 and 29.5% in NDCG@10. With this work, it becomes clear that item popularity dynamics explain much of the prediction ability of the modern neural recommender systems, shedding light on what neural models are learning from data and opening a new direction for improvement in the future.
2. Related Work
Figure 1. Model Architecture of PrepRec
Sequential Recommendation: Sequential recommenders model user behavior as a sequence of interactions, and aim to predict the next item that a user will interact. Early sequential recommenders adopt Markov chains and model item-item transition probability (Rendle et al., 2010; Shani et al., 2005). Later, with the advance of neural networks, some basic neural architectures have been explored for sequential recommendation, such as convolutional neural networks (Tang and Wang, 2018; Tuan and Phuong, 2017). Recurrent neural networks (Hidasi et al., 2016) and its variants, such as GRU4Rec (Hidasi et al., 2015) and GRU4Rec+ (Tan et al., 2016), have also been widely used for sequential recommendation due to the sequential nature of RNN-based architectures.
Recently, with the success of attention mechanisms in modeling sequential data, e.g., text sequences, some works have explored using attention mechanisms for sequential recommendation (Li et al., 2017; Liu et al., 2018). Attention-based models, e.g., Transformers (Vaswani et al., 2017) and BERT (Devlin et al., 2018a) show state-of-the-art performance in many NLP tasks. Inspired by the success of Transformers and BERT, some works have explored using attention-based models for sequential recommendation (Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020). However, these works only focus on the regular sequential recommendation and do not consider the zero-shot sequential recommendation. Some works take the timestamps of each interaction into consideration. (Li et al., 2020) looks at the time interval between two consecutive interactions and proposes a time-aware attention mechanism for sequential recommendation. Others (Lv et al., 2019; Ying et al., 2018; Tan et al., 2021) separate interaction sequences and categorize them to show the long-term and short-term interests of users. Some work on temporal recommendation model the change in users’ preferences and item popularities (Koren, 2009; Wu et al., 2017; Zhang et al., 2014). However, these works do not consider the dynamics of item popularity statistics and are also not capable of zero-shot sequential recommendation.
Symbol Description
$P(a^{t}_{j})$ popularity percentile of item $v_{j}$ at time $t$ over coarser time period
$P(b^{t}_{j})$ popularity percentile of item $v_{j}$ at time $t$ over finer time period
$\mathcal{P}^{t}_{j}$ popularity dynamics of item $v_{j}$ over a coarser time period up to time $t$
$\mathcal{H}^{t}_{j}$ popularity dynamics of item $v_{j}$ over a finer time period up to time $t$
$\mathbf{e}^{t}_{j}$ popularity dynamics embedding of item $v_{j}$ at time $t$
${\bm{T}}$ relative time interval encoding matrix
${\bm{P}}$ positional encoding matrix
$\mathbf{q}_{u}$ user embedding, computed from § 4.1.4
$h$ number of attention heads
$d$ embedding dimension
Table 1. Notation Table
Transfer Learning in Recommendation: Cross-domain transfer learning has been explored in recommender systems. Some works in the cross-domain recommendation literature focus on leveraging the relatively information-rich domain to improve the recommendation performance on the data-sparse domain (Hu et al., 2018; Li and Tuzhilin, 2020; Man et al., 2017). However, most of these works assume user or item overlap (Zhu et al., 2021; Zhao et al., 2020; Hu et al., 2018; Li and Tuzhilin, 2020; Man et al., 2017) for effective knowledge transfer. Other cross-domain literature focuses on the cold-start problem. There have been works proposed to tackle the cold-start user case (Lu et al., 2020; Zhu et al., 2021; Yuan et al., 2020; Feng et al., 2021) or the cold-start item case (Volkovs et al., 2017; Wei et al., 2021; Du et al., 2020; Liu et al., 2021). However, these works assume auxiliary information is available and can only handle one-sided cold-start problems (i.e., either cold-start users or cold-start items). Some other past works explore meta-learning on the cold-start problem and can handle both cold-start users and cold-start items (Lee et al., 2019; Dong et al., 2020) with application-dependent side information and fine-tuning on the target domain. Note that in this work, we are not trying to address the cold-start problem. Instead, we focus on the zero-shot transfer problem, where no auxiliary information is available in both domains.
Recently, pre-trained recommenders capable of zero-shot transfer have caught some attention in the community. (Ding et al., 2021; Hou et al., 2022) propose pre-trained models for sequential recommendations. However, they rely heavily on dataset-dependent auxiliary information (i.e., description text of items). (Wang et al., 2023) investigates the joint and marginal activity distribution of users and items, but are not suitable for the sequential recommendation task.
To summarize, prior works on sequential recommendation focus on learning high-quality representations for each item in the training set and are not generalizable across domains. Pre-trained sequential recommenders are capable of zero-shot recommendation to a new domain but rely heavily on auxiliary information of items, which is not always available in practice. In this paper, we propose a novel pre-trained sequential recommendation framework based on item popularity dynamics that are universal across datasets, and it is capable of zero-shot transfer without any auxiliary information.
3. Problem Definition
In this section, we formally define the research problems this paper addresses (i.e., regular sequential recommendation and zero-shot sequential recommendation) and introduce the notations used in this paper.
In this paper, we consider building the zero-shot sequential recommender system with only implicit feedback, i.e., no auxiliary information. In sequential recommendation, denote ${\bm{M}}$ as the implicit feedback matrix, $\mathcal{U}$ = $\{u_{1},u_{2},...,u_{|\mathcal{U}|}\}$ as the set of users, $\mathcal{V}$ = $\{v_{1},v_{2},...,v_{|\mathcal{V}|}\}$ as the set of items. The goal of sequential recommendation is to learn a scoring function, that predicts the next item $v_{u,t+1}$ given a user $u$ ’s interaction history $\{v_{u,1},v_{u,2},...,v_{u,t}\}$. Note that in this paper, since we model time explicitly, we assume access to the timestamp of each interaction, including the next item interaction. We argue that this is a reasonable assumption since the timestamp of the next interaction is always available in practice. For example, if Alice logs in to Netflix, Netflix will always know when Alice logs in and can predict the next movie for Alice. Formally, we define the scoring function as $\mathcal{F}(v^{t}|\mathcal{S}_{u},{\bm{M}})$, where $t$ is the time of the prediction.
Zero-shot Sequential Recommendation: In zero-shot sequential recommendation, given an interaction matrix ${\bm{M}}^{\prime}$ over $\mathcal{U}^{\prime}$ and $\mathcal{V^{\prime}}$, the goal is to produce a scoring function $\mathcal{F}^{\prime}$ without training on ${\bm{M}}^{\prime}$. In other words, the scoring function $\mathcal{F}^{\prime}$ has to be trained on a different interaction matrix ${\bm{M}}$ and then applied to ${\bm{M}}^{\prime}$ directly to predict the next item that user $u$ will interact with. In this paper, we study the zero-shot sequential recommendation that not only ${\bm{M}}$ and ${\bm{M}}^{\prime}$ are from different domains, but also both $\mathcal{U}\cap\mathcal{U}^{\prime}=\emptyset$ and $\mathcal{V}\cap\mathcal{V}^{\prime}=\emptyset$. In addition, we assume no auxiliary information is available in both domains.
Compared to previous works on zero-shot sequential recommendation, which assume access to side information, we pose a more challenging problem setting. We argue that this experimental setting is more universal and practical, and requires the model to learn the truly universal properties of the data.
4. PrepRec Framework
Here, we first introduce the model architecture of PrepRec (§ 4.1) and the training procedure (§ 4.2). Then, we formally define the zero-shot inference process (§ 4.3).
4.1. Model Architecture
First, we transform the user sequence $\{v_{u,1},v_{u,2},...,v_{u,|\mathcal{S}_{u}|}\}$ for each user $u$ into a fixed-length sequence $\mathcal{S}_{u}$ = $\{v_{u,1},v_{u,2},...,v_{u,L}\}$ via truncating or padding, where $L$ is a pre-defined hyper-parameter controlling the maximum length of the sequence. Note that the truncating is done by removing the oldest items in the sequence.
4.1.1. Universal Item Representation encoder
Previous works in sequential and pre-trained sequential recommender systems learn to represent items either through learnable embeddings specific to the item IDs, or through the item’s textual information. However, item IDs are not generalizable across domains and textual information is not always available and is application-dependent. In this work, we propose a novel item representation encoder that learns to represent items through the changes in their popularity histories over different time periods (popularity dynamics).
Formally, given an $v_{j}$ that has interaction in time period $t$, denoted as $v_{j}^{t}$, we define two popularity representations for $v_{j}^{t}$: popularity $\mathbf{p}_{j}^{t}\in\mathbb{R}^{k}$ over a coarse time period (e.g., month) and popularity $\mathbf{h}_{j}^{t}\in\mathbb{R}^{k}$ over a fine time period (e.g., week). To calculate $\mathbf{p}_{j}^{t}$ and $\mathbf{h}_{j}^{t}$ , we first count the number of interactions of $v_{j}^{t}$ over the two time period, denoted as $a_{j}^{t}\in\mathbb{R}$ (coarse period number of interactions) and $b_{j}^{t}\in\mathbb{R}$ (fine period number of interactions). Specifically, we calculate them as:
(1) $a_{j}^{t}=\sum_{m=1}^{t}\gamma^{t-m}c_{a}(v_{j}^{m}),\quad b_{j}^{t}=c_{b}(v_{j}^{t})$
where $\gamma\in\mathbb{R}$ is the discount factor and $c_{a}(v_{j}^{m})$ is the number of interactions of $v_{j}$ over a coarse time period $m$. Similarly, $c_{b}(v_{j}^{t})$ denotes the number of interactions of $v_{j}$ over a fine time period $t$. Note that we do not have a discounting factor when computing $b_{j}^{t}$ since we want it to capture the current popularity information, whereas $a_{j}^{t}$ captures the cumulative popularity of an item. We then calculate $a_{j}^{t}$ and $b_{j}^{t}$’s percentiles relative to their corresponding coarser and finer popularity distributions over all items at time $t$ and denote this by $P(a_{j}^{t})\in\mathbb{R}$ and $P({b_{j}^{t}})\in\mathbb{R}$, respectively. Finally, we encode $P(a_{j}^{t})$ and $P({b_{j}^{t}})$ into $k$ dimension vector representations $\mathbf{p}_{j}^{t}$ and $\mathbf{h}_{j}^{t}$ respectively. The encoding details can be found in Appendix A.
We now define the popularity dynamics of $v_{j}$ at time $t$ over the coarse time period (long-term horizon) to be $\{\mathbf{p}_{j}^{1},\mathbf{p}_{j}^{2},...,\mathbf{p}_{j}^{t-1}\}$, and over the fine time period (short-term horizon) as $\{\mathbf{h}_{j}^{1},\mathbf{h}_{j}^{2},...,\mathbf{h}_{j}^{t-1}\}$. The reason we use $t-1$ is to constrain access to future interactions and prevent information leakage, i.e., we do not have access to the popularity statistics of $v_{j}$ at time $t$ if we are at time $t$. For example, say an interaction happens on the second Wednesday in February, we consider the coarser and finer time period up until the end of January and the end of the first week in February respectively. To limit computation, we consider window sizes $m,n$ for $\mathcal{P}$ and $\mathcal{H}$ respectively. Formally, we now define the coarse popularity dynamics of $v_{j}$ at time $t$ as $\mathcal{P}_{j}^{t}=\{\mathbf{p}_{j}^{t-m},\mathbf{p}_{j}^{t-m+1},...,\mathbf{p}_{j}^{t-1}\}$, and the fine popularity dynamics of $v_{j}$ at time $t$ as $\mathcal{H}_{j}^{t}=\{\mathbf{h}_{j}^{t-n},\mathbf{h}_{j}^{t-n+1},...,\mathbf{h}_{j}^{t-1}\}$.
We then compute the embedding of item $v_{j}$ at time $t$ via the universal item representation encoder, defined as a function $\mathcal{E}(\mathcal{P}_{j}^{t},\mathcal{H}_{j}^{t})$ that learns to encode the popularity dynamics $\mathcal{P}_{j}^{t}$ and $\mathcal{H}_{j}^{t}$ into a $d$ dimension vector representation $\mathbf{e}_{j}^{t}$. Specifically, we have:
[…]
4.1.2. Relative Time Interval Encoding
Since we explicitly consider time when modeling sequences, we propose a relative time interval encoding to encode the time intervals between consecutive interactions of a user. Formally, given an interaction sequence $\mathcal{S}_{u}=\{v_{u,1},v_{u,2},...,v_{u,L}\}$ of user $u$, we define the time interval between $v_{u,j}$ and $v_{u,j+1}$ as $t_{u,j}=t(v_{u,j+1})-t(v_{u,j})$, where $t(v_{u,j})$ is the time that user $u$ interacts with item $v_{u,j}$. We then rank the time intervals of user $u$. Define the rank of relative time interval of $t_{u,j}$ as $r_{u,j}=\text{rank}(t_{u,j})$. The relative time interval encoding of interval $t_{u,j}$ is then defined as ${\bm{T}}_{r_{u,j}}\in\mathbb{R}^{d}$, where ${\bm{T}}\in\mathbb{R}^{L\times d}$, following the same setup in (Vaswani et al., 2017), is a fixed sinusoidal encoding matrix defined as:
(3) ${\bm{T}}_{i,2j}=\sin(\frac{i}{L^{2j/d}}),\quad{\bm{T}}_{i,2j+1}=\cos(\frac{i}{L^{2j/d}})$
We also tried a learnable time interval encoding, but this didn’t perform as well as the sinusoidal encoding. We hypothesize that the sinusoidal encoding is more generalizable across domains and the learnable encoding is more prone to overfitting.
4.1.3. Positional Encoding
As we will see in § 4.1.4, the self-attention mechanism does not take the positions of the items into account. Therefore, following (Vaswani et al., 2017), we also inject a fixed positional encoding for each position in a user’s sequence. Denote the positional embedding of a position $l$ as ${\bm{P}}_{l}\in\mathbb{R}^{d}$, where ${\bm{P}}\in\mathbb{R}^{L\times d}$. We compute ${\bm{P}}$ using the same formula in Equation 3. Again, we also tried a learnable positional encoding as presented in (Kang and McAuley, 2018; Sun et al., 2019), but it yielded worse results.
4.1.4. Popularity Dynamics-Aware Transformer Layer
We follow previous works in sequential recommendation (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2020) and propose an extension to the self-attention mechanism for modeling the popularity dynamics of items in a sequence by incorporating model components introduced in § 4.1.1§ 4.1.2, and § 4.1.3.
Given a user sequence $\mathcal{S}_{u}=\{v_{u,1},v_{u,2},...,v_{u,L}\}$, we first compute in input matrix ${\bm{E}}_{u}$ as:
(4) ${\bm{E}}_{u}=\left[\begin{aligned} &\mathbf{e}^{t}_{u,1}+{\bm{T}}_{r_{u,1}}+{\bm{P}}_{1}\\ &\mathbf{e}^{t^{\prime}}_{u,2}+{\bm{T}}_{r_{u,2}}+{\bm{P}}_{2}\\ &\vdots\\ &\mathbf{e}^{t*}_{u,L}+{\bm{T}}_{r_{u,L}}+{\bm{P}}_{L}\end{aligned}\right]$
$\mathbf{e}^{t}_{u,1},\mathbf{e}^{t^{\prime}}_{u,2},...,\mathbf{e}^{t*}_{u,L}$ is computed from Equation 2, ${\bm{T}}_{r_{u,1}},{\bm{T}}_{r_{u,2}},...,{\bm{T}}_{r_{u,L}}$ and ${\bm{P}}_{1},{\bm{P}}_{2},...,{\bm{P}}_{L}$ are computed following the procedure in § 4.1.2 and § 4.1.3 respectively.
Multi-Head Self-Attention. We adopt a widely used multi-head self-attention mechanism (Vaswani et al., 2017), i.e., Transformers. Specifically, it consists of multiple multi-head self-attention layers (denoted as MHAttn($\cdot$)), and point-wise feed-forward networks (FFN($\cdot$)). The multi-head self-attention mechanism is defined as:
(5) $\begin{split}\mathbf{z}_{u}&=\text{MHAttn}({\bm{E}}_{u})\\ \text{MHAttn}({\bm{E}}_{u})&=\text{Concat}(\text{head}_{1},...,\text{head}_{h}){\bm{W}}^{O}\\ \text{head}_{i}&=\text{Attn}({\bm{E}}_{u}{\bm{W}}_{i}^{Q},{\bm{E}}_{u}{\bm{W}}_{i}^{K},{\bm{E}}_{u}{\bm{W}}_{i}^{V})\end{split}$
where ${\bm{E}}_{u}$ is the input matrix computed from Equation 4, $h$ is a tunable hyper-parameter indicating the number of attention heads, ${\bm{W}}_{i}^{Q},{\bm{W}}_{i}^{K},{\bm{W}}_{i}^{V}\in\mathbb{R}^{d\times d/h}$ are the learnable weight matrices, and $\mathbf{W}^{O}\in\mathbb{R}^{d\times d}$ is also a learnable weight matrix. Attn is the attention function and is formally defined as:
(6) $\begin{split}\text{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})&=\text{softmax}(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d/h}})\mathbf{V}\end{split}$
The scale factor $\sqrt{d/h}$ is used to avoid large values of the inner product, which can lead to numerical instability.
Causality: In sequential recommendation, the prediction of the $s+1$ item that a user will interact with should only depend on the first $s$ items that the user has interacted with in the past. However, the $s$-th output of the multi-head self-attention layer contains all the input information. Therefore, as in (Li et al., 2020; Kang and McAuley, 2018), we do not let the model attend to the future items by forbidding links between $Q_{i}$ and $K_{j}$ ($j>i$) in the attention function.
Point-Wise Feed-Forward Network: While the multi-head self-attention mechanism can capture the dependencies between items in a sequence and incorporate all previous items’ information, it is still inherently a linear model. Therefore, to add nonlinearity and interactions between different embedding dimensions, we follow previous works in sequential recommendation (Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020) and apply the same point-wise feed-forward network to the output of each multi-head self-attention layer. Formally, suppose the output of the multi-head self-attention layer is $\textbf{z}_{u}$, the point-wise feed-forward network is defined as:
(7) $\begin{split}\text{FFN}(\textbf{z}_{u})&=\text{ReLU}(\textbf{z}_{u}{\bm{W}}_{1}+\mathbf{b}_{1}){\bm{W}}_{2}+\mathbf{b}_{2}\end{split}$
where ${\bm{W}}_{1}\in\mathbb{R}^{d\times d}$ and ${\bm{W}}_{2}\in\mathbb{R}^{d\times d}$ are learnable weight matrices, and $\mathbf{b}_{1}\in\mathbb{R}^{d}$ and $\mathbf{b}_{2}\in\mathbb{R}^{d}$ are learnable bias vectors.
Stacking Layers: As shown in previous works (Kang and McAuley, 2018), stacking multiple multi-head self-attention layers and point-wise feed-forward networks can potentially lead to overfitting and instability during the training. Therefore, we follow previous works (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2020) and apply layer normalization (Ba et al., 2016) and residual connections to each multi-head self-attention layer and point-wise feed-forward network. Formally, we have:
(8) $g(\mathbf{x})=\mathbf{x}+\text{Dropout}(g(\text{LayerNorm}(\mathbf{x})))$
$g(\mathbf{x})$ is either the multi-head self-attention layer or the point-wise feed-forward network. Therefore, for every multi-head self-attention layer and point-wise feed-forward network, we first apply layer normalization to the input, then apply the multi-head self-attention layer or point-wise feed-forward network, and finally apply dropout and add the input $\mathbf{x}$ to the layer output. The LayerNorm function is defined as:
(9) $\text{LayerNorm}(\mathbf{x})=\alpha\odot\frac{\mathbf{x}-\mu}{\sqrt{\sigma^{2}+\epsilon}}+\mathbf{\beta}$
where $\odot$ denotes the element-wise product, $\mu$ and $\sigma$ are the mean and standard deviation of $\mathbf{x}$, $\alpha$ and $\mathbf{\beta}$ are learnable parameters, and $\epsilon$ is a small constant to avoid numerical instability.
4.1.5. Prediction
Given a sequence $\mathcal{S}_{u}$ of user $u$, we denote the embedding of user $u$ after going through § 4.1.1 § 4.1.2 § 4.1.3 § 4.1.4 as $\mathbf{q}_{u}$ (Figure 1). Suppose at time $t^{+}$, we want to compute user $u$’s preference of item $v_{j}$, we first compute the item representation $\mathbf{e}_{j}^{t^{+}}$ from § 4.1.1. Then, we predict the score as the innner product of $\mathbf{q}_{u}$ and $\mathbf{e}_{j}^{t^{+}}$, formally:
(10) $\hat{y}(v^{t}_{j}|\mathcal{S}_{u})=<\mathbf{q}_{u},\mathbf{e}_{j}^{t^{+}}>$
There is no information leakage in the prediction process, i.e., we do not assume access to the popularity statistics of $v_{j}$ at time $t^{+}$ if we are at time $t^{+}$.
4.2. Training Procedure
Now we show how to train the PrepRec model. Similar to (Kang and McAuley, 2018), we adopt the binary cross entropy loss as the objective function, formally:
(11) $\begin{split}\mathcal{L}=-\sum_{\mathcal{S}_{u}\in\mathcal{S}}\sum_{z\in[1,2,...,L-1]}&[\log\sigma(\hat{y}(v^{t}_{z+1}|\mathcal{S}_{u,:z}))\\ &+\sum_{j^{\prime}\notin\mathcal{S}_{u}}\log\sigma(1-\hat{y}(v^{t}_{j^{\prime}}|\mathcal{S}_{u,:z}))]\\ \end{split}$
where $\mathcal{S}_{u,:z}\coloneqq\{v_{u,1},v_{u,2},...,v_{u,z}\}$. $v^{t}_{z+1}$ represents the $z+1$-th item in the sequence that happened at time $t$. We use Adam (Kingma and Ba, 2014) as the optimizer and train the model end-to-end. Note that compared to previous sequential recommenders, we do not have any parameters specific to the item IDs. Essentially, we are only optimizing ${\bm{W}}_{p}$ and parameters related to the multi-head self-attention mechanism.
4.3. Zero-shot Inference
Our inference setup is defined in § 3. For zero-shot transfer, we first compute two levels of popularity dynamics on the target dataset. Then, we directly apply the model pre-trained on the source dataset to get the prediction.
Formally, suppose we are given a pre-trained model $\mathcal{F}$ trained on ${\bm{M}}$. $\mathcal{F}$ is essentially the scoring function learned from source domain ${\bm{M}}$. Let us denote the interaction matrix of the target domain as ${\bm{M}}^{\prime}$. We first compute the popularity dynamics of each item in ${\bm{M}}^{\prime}$ over a coarser time period and a finer time period. Then, we apply the pre-trained model $\mathcal{F}$ to ${\bm{M}}^{\prime}$ and compute the prediction score as:
(12) $\hat{y}(v^{t}_{j^{\prime}}|\mathcal{S}^{\prime}_{u})=\mathcal{F}(v^{t}_{j^{\prime}}|\mathcal{S}^{\prime}_{u},{\bm{M}}^{\prime})$
Note that in this procedure, we use the pre-trained model $\mathcal{F}$ trained on domain ${\bm{M}}^{\prime}$ to predict the next item $v^{t}_{j^{\prime}}$ that user $u^{\prime}$ will interact with in domain ${\bm{M}}^{\prime}$. We do not use any auxiliary information in both domains. In addition, none of the parameters in $\mathcal{F}$ are updated during the inference process.
To summarize, in this section, we showed how to develop a pre-trained sequential recommender system based on the popularity dynamics of items. We enforce the structure of each interaction in the sequence by the positional encoding and introduce a relative time encoding for modeling time intervals between two consecutive interactions. In addition, we showed the training process and formally defined the zero-shot inference procedure. In the next section, we will evaluate PrepRec on multiple datasets.
5. Experiments
Dataset #users #items #actions avg length density
Office 101,133 27,500 0.74m 7.3 0.03%
Tool 240,464 73,153 1.96m 8.1 0.01%
Movie 70,404 40,210 11.55m 164.2 0.41%
Music 20,539 10,121 0.66m 32.2 0.32%
Epinions 30,989 20,382 0.54m 17.5 0.09%
Table 2. Dataset statistics
Dataset Office Tool Movie Music Epinions
Metric
R@10
N@10
R@10
N@10
R@10
N@10
R@10
N@10
R@10
N@10
General Recommender Systems
MostPop
0.450
0.272
0.459
0.274
0.586
0.361
0.519
0.327
0.438
0.296
BPR (Rendle et al., 2009)
0.457
0.289
0.363
0.216
0.747
0.477
0.646
0.434
0.568
0.397
NCF (He et al., 2017)
0.446
0.266
0.388
0.239
0.784
0.505
0.652
0.437
0.570
0.396
NGCF (Wang et al., 2019)
0.460
0.291
0.421
0.256
0.788
0.510
0.659
0.441
0.571
0.393
LightGCN (He et al., 2020)
0.465
0.293
0.463
0.275
0.793
0.512
0.665
0.447
0.575
0.396
Sequential Recommender Systems
Caser (Tang and Wang, 2018)
0.512
0.334
0.496
0.297
0.891
0.701
0.796
0.576
0.674
0.475
MARank (Yu et al., 2019)
0.488
0.312
0.468
0.282
0.888
0.683
0.739
0.501
0.628
0.436
SasRec (Kang and McAuley, 2018)
0.539
0.354
0.536
0.337
0.918
0.749
0.816
0.599
$0.705^{*}$
0.501
BERT4Rec (Sun et al., 2019)
$0.541^{*}$
$0.358^{*}$
$0.544^{*}$
$0.350^{*}$
0.900
0.728
$0.816^{*}$
$0.602^{*}$
0.702
$0.512^{*}$
TiSasRec (Li et al., 2020)
0.531
0.349
0.539
0.341
$0.918^{*}$
$0.752^{*}$
0.809
0.523
0.701
0.499
PrepRec
0.533
0.334
0.545
0.335
0.896
0.715
0.778
0.553
0.757
0.543
PrepRec $\Delta$
-1.5%
-7.1%
+0.2%
-4.4%
-2.5%
-4.9%
-4.6%
-8.1%
+7.3%
+6.1%
Interp 0.648 0.483 0.652 0.489 0.940 0.810 0.901 0.783 0.821 0.691
Interp $\Delta$
+19.8%
+34.9%
+19.9%
+39.7%
+2.4%
+7.7%
+10.4%
+30.1%
+16.5%
+35.0%
Table 3. Regular sequential recommendation results. We embolden the best results and mark the best baseline results with $*{}^{\prime}*^{\prime}$. Interp represents the interpolation results between PrepRec and BERT4Rec. PrepRec $\Delta$ denotes the performance difference between PrepRec and the best results among the selected baselines, similar for Interp $\Delta$. We want to emphasize that the regular sequential recommendation is not the main focus of this paper, rather, we include the results here to show the importance of modeling popularity dynamics in the sequential recommendation and establish a baseline of PrepRec’s performance. Despite not focusing on the performance of regular sequential recommendation, PrepRec achieves comparable performance to the state-of-the-art sequential recommenders, with only on average 0.2% worse than the best performing sequential recommenders in R@10 while having only a fraction of the model size (Table 9). After a simple post-hoc interpolation, we outperform the state-of-the-art sequential recommenders by 13.8% in R@10 on average.
We present extensive experiments on five real-world datasets to evaluate the performance of PrepRec, following the problem settings in § 3. We introduce the following research questions (RQ) to guide our experiments: (RQ1) How well can PrepRec perform under the regular, in-domain sequential recommendation setting? This helps us understand the importance of modeling popularity dynamics in sequential recommendation. (RQ2) Can PrepRec (popularity dynamics) generalize to zero-shot cross-domain and cross-application transfer? (RQ3) What are the factors that affect the performance of PrepRec ? In other words, how sensitive is PrepRec to different choices in constructing popularity dynamics? (RQ4) What is the effect of the amount of training data available for PrepRec ? This can be crucial if the pre-trained model is applied to a new domain with limited data.
5.1. Datasets and Preprocessing
We evaluate our proposed method on five real-world datasets, with varying applications, sizes, and density levels.
  • Amazon (Ni et al., 2019) is a series of product ratings datasets obtained from Amazon.com, split by product domain. We consider the Office and Tool product domains in our study.
  • Douban (Song et al., 2019) consists of three datasets across different domains, collected from Douban.com, a Chinese review website. We work with the Movie and Music datasets.
  • Epinions (Tang et al., 2012a, b) is a dataset crawled from product review site Epinions. We utilize the ratings dataset for our study.
The dataset statistics are presented in Table 2. We compute the density as the ratio of the number of interactions to the number of users times the number of items. The Tool dataset is the largest in terms of the number of users and items, with Office and Movie following, and Epinions and Music smallest. Both datasets from Douban (i.e., movie and music) are the densest, while the Amazon review datasets (i.e., office and tool) are the sparsest.
For fair evaluation, we follow the same preprocessing procedure as previous works (Kang and McAuley, 2018; Sun et al., 2019; Rendle et al., 2010; Tang and Wang, 2018), i.e., we binarize the explicit ratings to implicit feedback. In addition, for each user, we sort interactions by their timestamp and take these chronological sequences as input, using the second most recent action for validation, and the most recent action for testing. As described in § 4.1, our method directly utilizes interaction timestamps to discover long-term and short-term popularity trends. Note that existing sequential methods  (Sun et al., 2019; Kang and McAuley, 2018) use timestamps for sequence construction, so our method isn’t any more restrictive in application.
[…]
Metric
R@10
N@10
R@10
N@10
R@10
N@10
R@10
N@10
R@10
N@10
Reference: Regular Sequential Recommendation
MostPop
0.450
0.272
0.459
0.274
0.586
0.361
0.519
0.327
0.438
0.296
PrepRec
0.533
0.334
0.545
0.335
0.896
0.715
0.778
0.553
0.757
0.543
[…]
Office $\rightarrow$
0.560 0.346 0.875 0.665
0.731
0.531
0.711
0.523
Tool $\rightarrow$
0.548 0.336
0.866
0.648
0.752
0.542
0.713
0.511
[…]
0.497
0.296
0.508
0.302
0.773 0.551 0.757 0.545
[…]
0.485
0.304
0.496
0.312
0.739
0.469
0.723
0.513
[…]
0.446
0.296
0.470
0.302
0.783
0.588
0.719
0.540
[…]
5.2. Baselines and Experimental Setup
To assess the effectiveness of our proposed method, we compare against the following set of baselines. These baselines include classic general recommendation models (e.g., MostPop, BPR (Rendle et al., 2009), NCF (He et al., 2017), NGCF (Wang et al., 2019), LightGCN (He et al., 2020)) and state-of-the-art sequential recommendation models (e.g., Caser (Tang and Wang, 2018). MArank (Yu et al., 2019), SasRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019), TiSasRec (Li et al., 2020)). We provide more detailed descriptions of the baselines in Appendix B. To the best of our knowledge, we are the first to tackle the zero-shot sequential recommendation without any auxiliary information. Therefore, we do not include the existing pre-trained sequential recommenders (Hou et al., 2022; Ding et al., 2021) as they require textual descriptions of items, which we do not consider in the experiments.
Following previous works (Koren, 2008; He et al., 2017; Kang and McAuley, 2018; Sun et al., 2019), we adopt the leave-one-out evaluation method: for each user, we pair its test item with 100 unobserved items according to the user’s interaction history. Then we rank the test item for the user among the 101 total items. We use two standard evaluation metrics for top-$k$ recommendation: Recall@k (R@k) and Normalized Discounted Cumulative Gain@N (N@k). R@k measures the proportion of test items that appear in the top-$k$ recommendations, and N@k accounts for the ranking position of the test interaction. We report the average of R@k and N@k over all the test interactions.
We use publicly available Pytorch implementations for the baselines. For fair evaluation, we set dimension size $d$ to 50, max sequence length $L$ to $200$, and batch size to $128$ in all models. We use three layers (one base and two additional) in NCF and NGCF, and two attention layers for the sequential models (SasRec, BERT4Rec, TiSasRec, and ours). We use an Adam optimizer and tune the learning rate in the range $\{10^{-4},10^{-3},10^{-2}\}$ and set the weight decay to $10^{-5}$. We use the dropout regularization rate of $0.3$ for all models. We set $\gamma=0.5$ in Equation 1, whose reason we will discuss the reason in  § 5.5.1. We define the coarse and fine time period to be a month and a week respectively, and we fix the window size to be $m=12$ and $n=4$ for all datasets (§ 4.1.1). We train PrepRec for a maximum of 80 epochs. All experiments are conducted on a Tesla V100 using PyTorch. We repeat each experiment 5 times with different random seeds and report the average performance.
5.3. Regular Sequential Recommendation (RQ1)
We investigate the performance of PrepRec against the selected baselines. Note that the purpose of this paper is not to outperform the state-of-the-art sequential recommenders. Instead, we include the results here to investigate how well the popularity dynamics can capture users’ preferences in the sequential recommendation and can explain the state-of-the-art models. To no surprise, the selected sequential recommendation baselines are the best performant. However, PrepRec still achieves competitive performance- within 5% in R@10, - with the state-of-the-art sequential recommenders. On Epinions, PrepRec even outperforms all the sequential recommenders by 7.3%. We argue that this is particularly impressive given that PrepRec has significantly fewer model parameters (Table 9) and is capable of zero-shot transfer, with no side information.
PrepRec learns item representations through popularity dynamics, which is conceptually different from learning representations specific to each item ID. Therefore, we propose a simple post-hoc interpolation to investigate if modeling popularity dynamics can boost the performance of state-of-the-art sequential recommenders. We interpolate the scores from PrepRec with the scores from BERT4Rec as follows: $\hat{y}_{intp}(v^{t}_{j}|\mathcal{S}_{u})=\alpha*\hat{y}_{O}(v^{t}_{j}|\mathcal{S}_{u})+(1-\alpha)*\hat{y}_{S}(v_{j}|\mathcal{S}_{u})$ where $\hat{y}_{O}(v^{t}_{j}|\mathcal{S}_{u})$ and $\hat{y}_{S}(v_{j}|\mathcal{S}_{u})$ are the scores from PrepRec (Equation 10) and BERT4Rec, respectively. We then proceed to rank with the final interpolated scores. We set $\alpha=0.5$ for all datasets. After interpolation, there is a significant performance boost by up to 39.7% in N@10. This suggests that our model effectively captures dynamic item properties that are missing with the static item embeddings of SOTA sequential models. Gains are largest in the medium and low-density datasets (Epinions, Amazon), indicating that our model is particularly complimentary to existing methods in sparse datasets where item embeddings are less informative.
5.4. Cross-domain Zero-shot Transfer(RQ2)
For the zero-shot transfer experiments, we train PrepRec on the source dataset. We follow the zero-shot inference setting introduced in § 4.3. We make predictions using Equation 12 and report the results in Table 4. We also include the results of the regular sequential recommendation for reference. In the zero-shot setting, after training on source datasets, in the target dataset, PrepRec shows minimal performance reduction in most cases (i.e., 2.3% maximum reduction) and even outperforms PrepRec and selected sequential recommendation baselines that are trained on the target datasets. In particular, we test all combinations of source and target pairs. We find the zero-shot transfer result from Tool to Office (i.e., Tool is the source dataset while office is the target dataset) outperforms all the state-of-the-art sequential recommenders trained on office, i.e., 0.548 vs 0.541 (BERT4Rec) in R@10. Similarly, when the target is Music and Epinions, using Movie as the source dataset attains nearly identical performance with the reference models. We conjecture that source datasets with similar density levels to the target dataset generally achieve excellent generalization performance, indicating that our model captures relevant characteristics across the space of datasets. Overall, these results show our model’s effectiveness in transferring across domains and applications, performing at or above par with SOTA models trained on the target dataset, without any training on interaction data or side information. We provide a more detailed analysis (case study) in Appendix F.
5.5. Ablation Study (RQ3)
Dataset Music Office Epinions
Metric
R@10
N@10
R@10
N@10
R@10
N@10
$\gamma=0.5^{*}$ (weighted -pop)
0.778
0.553
0.533
0.334
0.757
0.543
$\gamma=0$ (Curr-pop)
0.749
0.542
0.512
0.328
0.689
0.496
$\gamma=1$ (cumul-pop)
0.695
0.452
0.539
0.330
0.733
0.505
Table 5. Recommendation results for varying the discounting factor $\gamma$ in § 4.1.1. $\gamma=0.5$ is the default setting, denoted by $*{}^{\prime}*^{\prime}$. We find that $\gamma=0.5$ generally outperforms the other two settings
5.5.1. Effect of discounting factor $\gamma_{a}$
We examine the effect of different preprocessing weights $\gamma$ used in popularity percentile calculation. In particular, $\gamma=1$ corresponds to the cumulative popularity, or in other words, at a given time period $t$, the overall number of interactions up to period $t$. On the other hand, $\gamma=0$ corresponds to the current popularity, or percentiles are calculated over interactions just in $t$, same as $b_{j}^{t}$ in Equation 1. When $\gamma=0.5$, it can be interpreted as interactions being exponentially weighted by time, with a half-life of 1 time period. We find that $\gamma=0.5$ generally outperforms the other two settings, with the largest gains of around 12% R@10 and 22% N@10 over cumul-pop in the dense Music dataset, and the largest gains over curr-pop in the sparser Office (4% N@10) and Epinions ($9\%$ R@10 and N@10) datasets. We suspect this is due to cumulative measures in denser datasets failing to capture recent trends due to the large historical presence, while current-only measures in sparser datasets convey too little or noisy information and lose the information of long-term trends. curr-pop shows decent performance on the Music dataset, suggesting that Music trends might be more cyclical and thus the current popularity is more informative.
Dataset Music Office Epinions
Metric
R@10
N@10
R@10
N@10
R@10
N@10
0.778
0.553
0.533
0.334
0.757
0.543
t-2
0.775
0.553
0.540
0.338
0.728
0.516
t-3
0.780
0.543
0.539
0.337
0.723
0.514
t-6
0.730
0.497
0.523
0.326
0.682
0.473
t-12
0.705
0.462
0.507
0.313
0.645
0.445
Table 6. Recommendation results for varying prediction time. Where $t-n$ denotes $n$ finer time periods ago. $t-1$ is the default setting, denoted by $*{}^{\prime}*^{\prime}$. The performance drops as $n$ goes larger, indicating the importance of recent trends in sequential recommendation.
5.5.2. Effect of Prediction Time
In PrepRec, we use the popularity trends up to some time as the context to construct universal item representations. To prevent information leakage, we can only use the popularity dynamics computed prior to the time of the interaction, as illustrated in § 4.1.1 and § 4.1.5. A natural choice would be to compute popularity dynamics as close to the actual interaction time as possible. However, experimenting with this can reveal the importance of recent trends in model performance Table 6. Thus, denoting $t-n$ to be $n$ fine time periods ago. We round the long-term horizon according to the change in $t$. Per expectation, we find that $t-6$ and $t-12$ drop performance significantly across all datasets, revealing that for the experimented datasets, capturing trends of the 6 recent weeks is crucial. Interestingly, we find that $t-2$ and $t-3$ perform worse in Epinions, but similarly to $t-1$ in Office and Music. We posit this could be due to dataset-specific cylic patterns where Music and Office may have more short-term recurring trends than Epinions, so not as much information is lost in the previous 2 weeks. This opens up possibilities for other axes which can be used to characterized recommender datasets, independent of size, sparsity, or overall duration.
5.6. Effect of Training Data Available(RQ4)
We investigate the influence of data sparsity on the performance of PrepRec and state-of-the-art sequential recommenders. While varying the training data size, we keep the test data fixed. We remove the oldest interactions from the training set and keep at least 2 training interactions per user to ensure the model is successfully trained.
We can see that the baseline sequential recommenders are sensitive to data sparsity, with significant performance drops when there is less training data. This is expected as the sequential recommenders learn item representations specific to item IDs, and the representation quality is highly dependent on the number of interactions. On the other hand, PrepRec learns item representation through popularity dynamics, showing that popularity dynamics are more robust to data sparsity. We argue that this is crucial for real-world applications where we want to transfer the model to a new domain with limited data. PrepRec is capable of learning from the limited data and outperforms the baseline sequential recommenders.
Figure 2. Recommendation results for varying the number of training interactions. 100% Training percentage corresponds to the result in Table 3. We decrease the training percentage by removing the oldest interactions. PrepRec is robust to data sparsity and outperforms all baseline sequential recommenders when fewer interactions are available.
5.7. Discussion
PrepRec shows excellent potential as a pre-trained sequential recommendation model, with competitive performance to the best baseline with a fraction of the model size and is capable of zero-shot transfer with minimal performance loss. We argue that PrepRec is particularly useful in the following scenarios: 1) initial sequential model when the data in the domain is sparse; 2) backbone for developing more complex sequential recommenders (i.e., prediction interpolation) 3) online recommendation settings. PrepRec does not explicitly model Item IDs, it can also easily generalize to new items/users with no retraining by just updating the popularity dynamics. We leave this as future work.
We also notice that the training process of state-of-the-art sequential recommenders assumes access to future interactions. For example, Alice has interacted with items $A,B,C$ in the past. The sequential recommenders will use $A,B,C$ to predict the next item $D$. If there’s another person, Bob, whose interaction history is $A,B,C,D$ and all of them happen after Alice’s test interactions $D$, the current sequential recommenders will use Bob’s interaction history $A,B,C,D$ to help Alice predict $D$. This is not realistic in the online recommendation setting. We found out that on average 31% of the training interactions are after the test interactions of the same dataset. More detailed statistics can be found in Appendix E. We argue that this is a major limitation of the current sequential recommenders and should be addressed in the future.
6. Conclusion
Inspired by the impact of pre-trained models, we propose PrepRec , a pre-trained sequential recommendation model that is capable of cross-domain zero-shot transfer without application-dependent side information. PrepRec learns to represent items based on their popularity dynamics and computes user representations via a popularity dynamics-aware transformer layer. Through extensive experiments, we empirically show that PrepRec can achieve comparable performance to or even outperform state-of-the-art sequential recommenders on regular sequential recommendation tasks. PrepRec also shows excellent generalizability and can zero-shot transfer to a new domain with minimal performance reduction, indicating that popularity dynamics play an important role in sequential recommendation and are crucial for developing generalizable sequential recommenders.
Appendix A Popularity Encoding
We provide details for the item popularity encoding technique introduced in Section 4.1.1. Denote the popularity encoder as $E_{p}$, which takes in a percentile $P(x)\in\mathbb{R}^{+}$. Suppose given the popularity percentile $P(a_{j}^{t})\in\mathbb{R}^{+}$ over a coarse time period $t$, the coarse level popularity vector representation $\mathbf{p}_{j}^{t}\in\mathbb{R}^{k}$ and is computed as follows:
$\mathbf{p}_{j}^{t}$ $=E_{p}(P(a_{j}^{t}))$
$(\mathbf{p}_{j}^{t})_{i}$ =$=\begin{cases}1-\{\frac{P}{10}\},&\text{if }i=\lfloor\frac{P}{10}\rfloor\\ \{\frac{P}{10}\},&\text{if }i=\lfloor\frac{P}{10}\rfloor+1\\ 0,&\text{otherwise}\end{cases}$
[…]
Appendix B Baselines Description
In this section, we provide a brief description of the baselines used in our experiments. We tune the hyper-parameters for each baseline on the validation set and report the best results on the test set.
  • MostPop is a non-personalized baseline that recommends the most popular items in the training set.
  • BPR (Rendle et al., 2009) is a learning-to-rank method that learns latent representations for users and items through matrix factorization.
  • NCF (He et al., 2017) learns user and item embeddings through a multi-layer perceptron instead of matrix factorization.
  • NGCF (Wang et al., 2019) uses a bipartite graph to model user-item interactions through a graph convolutional network (Kipf and Welling, 2016).
  • LightGCN (He et al., 2020) is a simplified version of NGCF that removes the feature transformation and nonlinear activation.
  • Caser (Tang and Wang, 2018) learns convolutional filters on user interaction sequences over the sequence order and latent dimensions.
  • MArank (Yu et al., 2019) captures individual and union-level item dependency through multi-order attention and a residual network.
  • SasRec (Kang and McAuley, 2018) uses self-attention to model user embeddings through its sequence of latent item representations.
  • BERT4Rec (Sun et al., 2019) improves on SasRec by using a bi-directional transformer to capture interactions from both sides.
  • TiSasRec (Li et al., 2020) additionally embeds the raw time interval between consecutive item interactions in the sequence.
Appendix C Additional Experimental Results
Figure 3. Recommendation results for varying user activity levels. Subsequent groups along the $x$-axis have higher activity levels, i.e., user group 1 corresponds to the least active group of users while group 5 corresponds to the most active ones. We find that PrepRec performs relatively better for users with more interactions
C.1. Performance by User Group
In this experiment, we study the performance of our model on different user activity levels (Figure 3). We divide users into 5 evenly separated groups based on their number of interactions and present the NDCG@10 for each group. We can see from the figure that the gap between baselines and PrepRec decreases as user activity increases for Office and Movie, while PrepRec outperforms baselines for all user groups on Epinions with significant margins. This suggests that in the presence of many interactions, popularity dynamics can capture much of the user’s preferences without using item IDs.
C.2. Parameter Sensitivity
C.2.1. Effect of Window Size
Dataset Music Office Epinions
Metric
R@10
N@10
R@10
N@10
R@10
N@10
36m
0.788
0.529
0.515
0.332
0.751
0.556
24m
0.790
0.553
0.530
0.333
0.748
0.542
0.778
0.553
0.533
0.334
0.757
0.543
6m
0.762
0.546
0.540
0.340
0.761
0.542
3m
0.766
0.545
0.520
0.331
0.764
0.530
2m
0.768
0.528
0.511
0.325
0.745
0.526
1m
0.784
0.543
0.510
0.325
0.747
0.527
Table 7. Recommendation results for varying coarse level popularity dynamics window size $m$ (§ 4.1.1).
We assess the effect of different numbers of window size $m$ used in constructing the coarse level popularity dynamics introduced in § 4.1.1 (Table 7). For all our datasets, we fix $n=4$ for the finer horizon, but experiment with different $m$ for the coarser horizon. The construction method for popularity dynamics and the model’s item embedding sizes remain constant, so we’re effectively trading off incorporating more historical information at the risk of losing nuances in recent trends. Results are presented in Table 7. We find that 6 and 12-month windows are generally best for Office and Epinions, although NDCG@10 improves close to monotonically with window size for Epinions. There isn’t a clear trend for Music, with the 24-month windows performing best, but the 1-month window is competitive as well. This suggests that denser datasets may have more cyclical popularity trends that are captured equally well with shorter and longer windows, while sparser datasets are more sensitive and require more fine-tuning for optimal window size.
Figure 4. Recommendation results for varying the preprocessing discounting factor. $\gamma=0.25$ and $\gamma=0.5$ yield competitive results over the various datasets, while $\gamma=0.75$ is slightly worse.
C.2.2. Effect of Discounting Factor
[…]
C.3. Time and Positional Encoding Ablations
Our model architecture uses a relative time embedding by encoding the relative magnitude time interval between consecutive interactions. Here we assess the importance of this time embedding by comparing the full model, denoted as PrepRec, to a model without it, denoted w/o ${\bm{T}}$. We found that with the addition of relative time encoding ${\bm{T}}$, the performance generally improves. In addition, previous works (Sun et al., 2019; Kang and McAuley, 2018) consider learnable as opposed to fixed positional embeddings. We experiment with learnable positional encodings, denoted Learnable ${\bm{P}}$, for the regular, in-domain setting. We also note that learnable embeddings may be prone to overfitting, and thus won’t generalize well to new domains in zero-shot transfer. We reason that fixed embeddings achieve better performance due to the smaller number of overall parameters our model has (Table 9), resulting in strong test performances for simpler models.
Dataset Music Office
PrepRec N@10
0.553
0.334
w/o ${\bm{T}}$ N@10
0.514
0.334
Learnable ${\bm{P}}$ N@10
0.546
0.330
Table 8. Ablation study for time and positional embeddings. We find that the relative time embedding is important for recommendation performance. On the otherhand, we find that learnable positional embeddings do not perform as well as fixed embeddings.
Appendix D Model Size Analysis
Dataset Office Tool Movie Music Epinions
SasRec
1.331m
3.581m
2.044m
0.542m
1.054m
BERT4Rec
2.687m
7.233m
4.126m
1.094m
2.127m
TiSasRec
1.367m
3.617m
2.127m
0.578m
1.09m
PrepRec
0.045m
0.045m
0.045m
0.045m
0.045m
Table 9. Comparison of model sizes (i.e., number of learnable parameters) over different datasets. PrepRec uses significantly fewer number of parameters than baselines, ranging from 12 to 90x smaller.
In this section, we compare model complexities over different datasets. All the baseline methods learn item embeddings specific to their explicit item IDs, which scale with the number of items in the dataset. This can become a bottleneck for large-scale datasets and prevents zero-shot generalization to new datasets. In contrast, PrepRec learns item representations by modeling their popularity dynamics, which are independent of the number of items and will not scale with the number of items. We compare the model sizes of PrepRec and sequential recommendation baselines in Table 9.
We find that PrepRec uses significantly fewer learnable parameters than sequential recommendation baselines. Indeed, we can see the model size remains constant across datasets of varying sizes, as the number of model parameters does not scale with the dataset, making PrepRec more suitable for large-scale online platforms. Despite its much smaller size, PrepRec is able to achieve competitive performance with baselines.
Appendix E Time Leakage Analysis
Dataset Office Tool Movie Music Epinions
# Future Actions
18.22
18.37
597.56
72.9
27.39
% Future Actions
30.46%
24.39%
45.96%
28.32%
28.70%
Table 10. Time leakage statistics. The first row measures the average number of future interactions of the associated test item included in the training set. The second row measures the average proportion of all of the associated item’s training set interactions that are future interactions.
Sequential recommendation baselines like SasRec, TiSasRec, and BERT4Rec are trained on the users’ sequences without considering the interaction time. As the models train over different user sequences, the item representation of a given item will be updated without any consideration of the exact time at which the interaction occurred. Concretely, since the validation and test sets are relative to each user, a given item $v_{j}$ that interacted with user $u_{i}$ at time $t$ and user $u_{i^{\prime}}$ at time $t^{\prime}$ ($t^{\prime}>t$). The item $v_{j}$ may be part of the training interactions for user $u_{i^{\prime}}$ but the testing interactions for user $u_{i}$. So upon inference time for user $u_{i}$, the item embedding for $v_{j}$ will include information about a future interaction with user $u_{i^{\prime}}$, which would not be available in practice. In contrast, since our method represents items by popularity embeddings up to the time of the interaction, there is no time leakage because there’s no associated item ID, and the training has been conducted only on the universal item representation based on popularity dynamics.
Group
1
2
3
4
5
Consecutive Interactions Time Deviation normal-$\to$
Reg R@10
0.757
0.751
0.781
0.785
0.813
ZS R@10
0.729
0.740
0.771
0.793
0.832
$\Delta$ R@10
-1.5%
-1.2%
+1.0%
Reg N@10
0.518
0.522
0.555
0.567
0.603
ZS N@10
0.4999
0.508
0.546
0.578
0.623
$\Delta$ N@10
-2.6%
-1.6%
+1.9%
First Interaction Time normal-$\to$
Reg R@10
0.795
0.782
0.757
0.762
0.791
ZS R@10
0.780
0.754
0.755
0.771
0.804
$\Delta$ R@10
-1.9%
-0.3%
1.1%
+1.6%
Reg N@10
0.568
0.559
0.541
0.535
0.560
ZS N@10
0.546
0.532
0.531
0.548
0.597
$\Delta$ N@10
-2.0%
2.4%
Last Interaction Time normal-$\to$
Reg R@10
0.656
0.780
0.816
0.750
0.885
ZS R@10
0.614
0.766
0.807
0.795
0.884
$\Delta$ R@10
-1.9%
-1.0%
-0.2%
Reg N@10
0.431
0.548
0.633
0.500
0.652
ZS N@10
0.402
0.529
0.598
0.547
0.679
$\Delta$ N@10
Table 11. Zero-shot Transfer Analysis. We compare the in-domain (trained on target) and zero-shot cross-domain (trained and source and transferred to target) performance of PrepRec. Reg stands for the model performance trained on the target dataset, while ZS stands for the zero-shot cross-domain performance. We use the music as the target dataset and the movie as the source dataset. We divide users into 5 groups based on their interaction time deviation, first interaction time, and last interaction time.
Appendix F Zero-Shot Transfer Analysis
The zero-shot transfer results in Table 4 show that the proposed method can generalize quite well from source to target datasets, achieving competitive performance with baselines and our model trained on the target dataset, despite having no access to the target dataset during training. In this section, we further analyze the performance of zero-shot transfer by comparing its performance to regularly trained PrepRec across different user groups. We consider the Movie dataset as the source and Music dataset as the target, and contrast our regular, in-domain model, denoted Reg with our zero-shot, cross-domain model, denoted ZS. All statistics increase for larger group numbers, as represented by $\to$. $\Delta$ denotes the difference in metrics between ZS and Reg. Statistical significance using the two-sided t-test for independence is indicated by $*{}^{*}$ for $p<0.05$ and ${}^{**}$ for $p<0.01$ in the difference rows.
We find statistically significant results for three different time-based metrics: the deviation in time between consecutive interactions, the time of the first interaction, and the time of the last interaction. Concretely, for each of these metrics, we sort all users by the metric and divide them into 5 groups, with Group 1 being the smallest metric values and Group 5 being the largest, and assess performance. For consecutive interaction deviation, Reg performs better on users with smaller deviations or more frequent interactions, while ZS performs better on users with larger deviations or more spread out interactions. We posit that due to the larger amount of training data, the Movie-trained ZS model captures longer dynamic trends better, but doesn’t cature some shorter-term, idiosyncratic patterns. Also users with earlier first interactions and earlier last interactions also perform better with Reg, while predictions for later first interactions and later last interactions are better with ZS. This could be due to items later in the dataset having more precise popularity data, leading to the pretrained ZS outperforming the in-domain Reg model.