### v2

# Pre-trained Sequential Recommendation Framework via Popularity Dynamics

Abstract.
Sequential recommenders models sequential users’ behavior and are the fundamental building blocks of many online applications, e.g., e-commerce, video streaming, and social media. While effective, they learn item embeddings one-on-one mapped to explicit item IDs, which are unique in different domains, making them hard to generalize to new recommendation domains, whereas pre-trained models have shown great success in generalizing to new application domains in peer AI fields. Pioneer works on pre-trained sequential recommenders rely on application-dependent auxiliary information and can only generalize to the related domains within the same application, e.g., online retail. In this work, we aim to set a baseline for pre-trianed sequential recommenders by tackling the most challenging task: zero-shot cross-domain and cross-application sequential recommendation without auxiliary information. We present a novel pre-trained sequential recommendation framework: PrepRec. We learn universal item and sequence representations by modeling item dynamics and propose a popularity dynamics-aware transformer architecture. Through extensive experiments on five real-world datasets, we show that PrepRec, without any auxiliary information, can zero-shot adapt to new application domains and achieve competitive performance compared to state-of-the-art sequential recommender models. In addition, with a simple post-hoc interpolation, PrepRec can improve the performance of existing sequential recommenders on average by 12.4% in Recall@10 and 22% in NDCG@10. We provide an anonymized implementation of PrepRec at https://anonymous.4open.science/r/PrepRec--2F60/.
Recommender System, Neural Collaborative Filtering, Sequential Recommendation, Zero-shot Sequential Recommendation
copyright: acmcopyrightjournalyear: 2018doi: 10.1145/1122445.1122456copyright: noneconference: Woodstock ’18: ACM Symposium on Neural Gaze Detection; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-$9$-9/18/06
1. Introduction
Modeling sequential user behavior is critical to the success of online applications such as e-commerce, video streaming, and social media. Despite essential innovations for tackling the sequential recommendation task (Rendle et al., 2010; Hidasi et al., 2015; Tan et al., 2016; Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020), these models have to be trained from scratch for each application domain because they learn representations for each item (Sun et al., 2019; Kang and McAuley, 2018) in the domain. Further, to maintain high performance, we need to retrain the models when there is a large influx of new users and items. Since the models learn representations for each item, they cannot generalize to new domains. Thus, in this paper, we ask: can we build a pre-trained sequential recommender system capable of cross-domain and cross-application transfer without any auxiliary information? (e.g., using the model trained for online shopping in the US to predict the next movie that a user in India will watch).
At first glance, developing pre-trained sequential recommenders for cross-domain inference seems impossible. While we see pre-trained language (Devlin et al., 2018; Brown et al., 2020; OpenAI, 2023; Raffel et al., 2020; Liu et al., 2019) and vision models (Dosovitskiy et al., 2020; He et al., 2016; Radford et al., 2021) show excellent generalizability across datasets and applications, being able to achieve state-of-the-art performance by just a few fine-tuning steps (Devlin et al., 2018; Liu et al., 2019) or even without any training (OpenAI, 2023; Brown et al., 2020) (i.e., zero-shot transfer), there are essential differences. The representations learned by the pre-trained language model seem universal since the training domain and the application domain (e.g., text prediction and generation) share the same language and vocabulary, supporting the effective reuse of the word representations. However, in the cross-domain recommendation problem, the items are distinct across domains in sequential recommendation datasets (e.g., grocery items vs movies). Therefore, forming such generalizable correspondence is nearly impossible if we learn specific representations for each item within each domain. Recent work explores pre-trained models for sequential recommendation (Ding et al., 2021; Hou et al., 2022; Hao et al., 2021) within the same application (e.g., online retail). However, they assume access to metadata of items (e.g., item description), which is application-dependent and is often not generalizable to other applications. These models cannot learn universal representations of items; instead, they bypass the representation learning problem by using additional item-side information. This paper tackles the more challenging task of zero-shot, cross-domain and cross-application sequential recommendations without any auxiliary information.
Present Work: Our crucial insight is to learn the universal item representations based on the item popularity dynamics and use these representations for cross-domain and cross-application zero-shot prediction. We know that the marginal distribution of user and item activities is universal and heavy tailed across datasets, supported by prior work in network science (Barabási and Albert, 1999; Barabasi, 2005) and by experiments in recommender systems (Salganik et al., 2006). In addition, recent work in recommender systems suggests that the popularity dynamics of items are also crucial for predicting users’ behaviors (Ji et al., 2020). Therefore, in this paper, we explicitly model the popularity dynamics of items and propose a novel pre-trained sequential recommendation framework: PrepRec. We learn universal item representations based on their popularity dynamics, instead of their explicit item IDs or auxiliary information. We encode the relative time interval between two consecutive interactions via relative-time encoding and the position of each interaction in the sequence by positional encoding. Using physical time ensures that the predictions are not anti-causal, i.e., using the future interactions to predict the present. We propose a popularity dynamics-aware transformer architecture for learning universal sequence representations. We show that it is possible to build a pre-trained sequential recommender system capable of cross-domain and cross-application transfer without any auxiliary information. Our key contributions are as follows:
Universal item and sequence representations::
We are the first to learn universal item and sequence representations for sequential recommendation without any auxiliary information by exploiting item popularity dynamics. In contrast, prior research learns item representations for each item ID or through item auxiliary information. We learn universal item representations by modeling item popularity dynamics of two temporal resolutions: coarse and fine-grained. We learn universal sequence representations using a carefully designed popularity dynamics-aware transformer architecture. These universal item and sequence representations make possible pre-trained sequential recommender systems capable of cross-domain and cross-application transfer without any auxiliary information.
Zero-shot transfer without auxiliary information::
We propose a new challenging setting for pre-trained sequential recommender systems: zero-shot cross-domain and cross-application transfer without any auxiliary information. In contrast, previous pre-trained sequential recommenders requires overlapping users (Yuan et al., 2020), application-dependent auxiliary information (Ding et al., 2021; Hou et al., 2022; Hao et al., 2021; Hou et al., 2023; Wang et al., 2023b), and are few-shot adapted to related domains within the same application (Hou et al., 2022, 2023; Wang et al., 2023b). Assuming no user or item metadata posts a more challenging but crucial problem as we want to set a baseline for the performance of pre-trained sequential recommenders. Using metadata can only make the problem easier and improve the performance.
With extensive experiments, we empirically show that PrepRec has excellent generalizability across domains and applications. Remarkably, had we trained a state-of-the-art model from scratch for the target domain, instead of zero-shot transfer using PrepRec, the maximum performance gain would have been only 4%. In addition, we show that with a post-hoc interpolation, PrepRec can outperform the state-of-the-art sequential recommender system on average by 12.4% in Recall@10 and 22% in NDCG@10. We attribute the improvements to the performance gains over long-tail items, which we show in the qualitative analysis. With this work, we set a baseline for pre-trained sequential recommenders and show that it is possible to build a pre-trained sequential recommender system under the most challenging setting: cross-domain and cross-application transfer without any auxiliary information.
2. Related Work
Figure 1. Model Architecture of PrepRec
Sequential Recommendation: Sequential recommenders model user behavior as a sequence of interactions, and aim to predict the next item that a user will interact with. Early sequential recommenders adopt Markov chains and model item-item transition probability (Rendle et al., 2010; Shani et al., 2005). Later, with the advance of neural networks, some basic neural architectures have been explored for sequential recommenders, such as convolutional neural networks (Tang and Wang, 2018; Tuan and Phuong, 2017). Recurrent neural networks (Hidasi et al., 2016) and its variants, such as GRU4Rec (Hidasi et al., 2015) and GRU4Rec+ (Tan et al., 2016), have also been widely used for sequential recommendation due to the sequential nature of RNN-based architectures.
With the success of Transformer (Vaswani et al., 2017) in modeling sequential data, e.g., text sequences, some works adopt the transformer architecture for sequential recommendation (Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020). Additionally, (Li et al., 2020) considers the timestamps of each interaction and looks at the time interval between two consecutive interactions and proposes a time-aware attention mechanism. Some works (Lv et al., 2019; Ying et al., 2018; Tan et al., 2021) separate interaction sequences and categorize them to show the long-term and short-term interests of users. Temporal sequential recommenders (Koren, 2009; Wu et al., 2017; Zhang et al., 2014) models the change in users’ preferences. These works, while achieving state-of-the-art performance, only focus on the regular sequential recommendation and cannot transfer to other domains.
Cross-domain Recommendation: Cross-domain recommendation literature leverages the information-rich domain to improve the recommendation performance on the data-sparse domain (Hu et al., 2018; Li and Tuzhilin, 2020; Man et al., 2017). However, most of these works assume user or item overlap (Zhu et al., 2021; Zhao et al., 2020; Hu et al., 2018; Li and Tuzhilin, 2020; Man et al., 2017) for effective knowledge transfer. Other cross-domain literature focuses on the cold-start problem (Lu et al., 2020; Zhu et al., 2021; Feng et al., 2021; Volkovs et al., 2017; Wei et al., 2021; Du et al., 2020; Liu et al., 2021; Lee et al., 2019; Dong et al., 2020). Note that we are not trying to address the cold-start problem. Instead, we focus on building a pre-trained sequential recommendation framework that can zero-shot transfer to a new domain without any auxiliary information.
Pre-trained Sequential Recommenders: Recently, with the success of pre-trained language models(Devlin et al., 2018; Brown et al., 2020; OpenAI, 2023), pre-trained recommenders have caught the attention of the community. ZESRec (Ding et al., 2021) is capable of zero-shot sequential recommendations. However, it only works for closely related domains and relies on application-dependent auxiliary information (i.e., description text of items). PeterRec (Yuan et al., 2020) requires overlapping users in both domains. On the other hand, finetuning-based models, e.g., MISSRec (Wang et al., 2023b), UnisRec (Hou et al., 2022), and VQ-Rec (Hou et al., 2023), are not designed for zero-shot sequential recommendation and works within the same application (e-commerce) since they rely on application-dependent auxiliary information.  (Wang et al., 2023a) investigates the joint and marginal activity distribution of users and items, but are not suitable for the sequential recommendation task.
To summarize, prior works on sequential recommendation focus on learning high-quality representations for each item in the training set and are not generalizable across domains. Pre-trained sequential recommenders are evaluated on closely related domains and platforms and rely heavily on application-dependent auxiliary information of items.
3. Problem Definition
In this section, we formally define the research problems this paper addresses (i.e., regular sequential recommendation and zero-shot sequential recommendation) and introduce our notations.
In sequential recommendation, denote ${\bm{M}}$ as the implicit feedback matrix, $\mathcal{U}$ = $\{u_{1},u_{2},...,u_{|\mathcal{U}|}\}$ as the set of users, $\mathcal{V}$ = $\{v_{1},v_{2},...,v_{|\mathcal{V}|}\}$ as the set of items. The goal of sequential recommendation is to learn a scoring function, that predicts the next item $v_{u,t}$ given a user $u$ ’s interaction history $\mathcal{S}_{u}=\{v_{u,1},v_{u,2},...,v_{u,t-1}\}$. Note that in this paper, since we model time explicitly, we assume access to the timestamp of each interaction, including the next item interaction. We argue that this is a reasonable assumption since the timestamp of the next interaction is always available in practice. For example, if Alice logs in to Netflix, Netflix will always know when Alice logs in and can predict the next movie for Alice. Formally, we define the scoring function as $\mathcal{F}(v^{t}|\mathcal{S}_{u},{\bm{M}})$, where $t$ is the time of the prediction.
Zero-shot Sequential Recommendation: Given two domains ${\bm{M}}$ and ${\bm{M}}^{\prime}$ over $\mathcal{U},\mathcal{V}$ and $\mathcal{U}^{\prime},\mathcal{V^{\prime}}$ respectively, we study the zero-shot recommendation problem in the scenario where the domains are different (${\bm{M}}\cap{\bm{M}}^{\prime}=\varnothing$), users are disjoint ($\mathcal{U}\cap\mathcal{U^{\prime}}=\varnothing$), and item sets are unique ($\mathcal{V}\cap\mathcal{V^{\prime}}=\varnothing$). The goal is to produce a scoring function $\mathcal{F}^{\prime}$ without training on ${\bm{M}}^{\prime}$ directly. In other words, the scoring function $\mathcal{F}^{\prime}$ has to be trained on a different interaction matrix ${\bm{M}}$. Furthermore, we assume there is no metadata associated with users or items, which makes the problem particularly challenging but crucial to study. We want to set a baseline for pre-trained sequential recommenders, using metadata can only improve and simply the problem.
4. PrepRec Framework
We first introduce the model architecture of PrepRec (§ 4.1) then the training procedure (§ 4.2). Finally, we formally define the zero-shot inference process (§ 4.3).
4.1. Model Architecture
The first step of building a pre-trained sequential recommender is to learn universal item representations. Prior works in sequential and pre-trained sequential recommender systems learn to represent items either through learnable embeddings specific to their item IDs or through their textual metadata, largely limiting the generalizability of the model under various scenarios, e.g., different domains, platforms, and applications. Our solution is to exploit the item popularity statistics available in every domain to learn universal item representations. Instead of explicit item IDs or auxiliary information, we learn to represent items at a given timestamp through the changes in their popularity histories over different periods (popularity dynamics). We propose a popularity dynamics-aware Transformer architecture that obtains the representation of users’ behavior sequences through item popularity dynamics.
4.1.1. Item Popularity Encoder
We learn to represent items based on their popularity dynamics, i.e., changes in their popularity histories. Intuitively, we should model items’ popularity by two horizons: long-term and short-term. Long-term horizons reflect the overall popularity of items, whereas short-term horizons should capture the recent trends in the domain. For example, the long-term popularity of a winter coat measures how popular is the coat in general, while its short-term popularity reflects more temporal changes, e.g., season, weather conditions, and fashion trends. Therefore, consider an item $v_{j}$ that has interaction at time $t$, denoted as $v_{j}^{t}$, we define two popularity representations for $v_{j}^{t}$: popularity $\mathbf{p}_{j}^{t}\in\mathbb{R}^{k}$ over a coarse period (e.g., month) and popularity $\mathbf{h}_{j}^{t}\in\mathbb{R}^{k}$ over a fine period (e.g., week).
To calculate $\mathbf{p}_{j}^{t}$ and $\mathbf{h}_{j}^{t}$, we first calculate the popularities of $v_{j}^{t}$ over the two horizons, denoted as $a_{j}^{t}\in\mathbb{R}^{+}$ (coarse period number of interactions) and $b_{j}^{t}\in\mathbb{R}^{+}$ (fine period number of interactions). Specifically, we calculate them as:
(1) $a_{j}^{t}=\sum_{m=1}^{t}\gamma^{t-m}c_{a}(v_{j}^{m}),\quad b_{j}^{t}=c_{b}(v_{j}^{t})$
where $\gamma\in\mathbb{R}^{+}$ is a pre-defined discount factor and $c_{a}(v_{j}^{m})$ is the number of interactions of $v_{j}$ over a coarse time period $m$. Similarly, $c_{b}(v_{j}^{t})$ denotes the number of interactions of $v_{j}$ over a fine period $t$. We do not impose the discounting factor when computing $b_{j}^{t}$ since we want it to capture the current popularity information, whereas $a_{j}^{t}$ captures the cumulative popularity of an item over a longer horizon.
The popularity of an item is a real value, which is not comparable across items from different domains. Therefore, we calculate the percentiles of $a_{j}^{t}$ and $b_{j}^{t}$ relative to their corresponding coarser and finer popularity distributions over all items at time $t$, denoted as $P(a_{j}^{t})\in\mathbb{R}^{+}$ and $P({b_{j}^{t}})\in\mathbb{R}^{+}$, respectively.
We now encode the popularity percentiles $P(a_{j}^{t})$ and $P({b_{j}^{t}})$ into $k$ dimensional vector representations $\mathbf{p}_{j}^{t}$ and $\mathbf{h}_{j}^{t}$ respectively. Denote the popularity encoder as $E_{p}:\mathbb{R}^{+}\rightarrow\mathbb{R}^{k}$, which takes in a percentile value. Suppose given the popularity percentile $P(a_{j}^{t})\in\mathbb{R}^{+}$ over a coarse time period $t$, the coarse level popularity vector representation $\mathbf{p}_{j}^{t}\in\mathbb{R}^{k}$ is computed as follows:
$\mathbf{p}_{j}^{t}$ $=E_{p}(P(a_{j}^{t}))$
$(\mathbf{p}_{j}^{t})_{i}$ =$=\begin{cases}1-\{\frac{P}{k-1}\},&\text{if }i=\lfloor\frac{P}{k-1}\rfloor\\ \{\frac{P}{k-1}\},&\text{if }i=\lfloor\frac{P}{k-1}\rfloor+1\\ 0,&\text{otherwise}\end{cases}$
[…]
4.1.2. Universal Item Representation
We now define the popularity dynamics of $v_{j}$ at time $t$ over the coarse period (long-term horizon) to be $\mathcal{P}_{j}^{t}=\{\mathbf{p}_{j}^{1},\mathbf{p}_{j}^{2},...,\mathbf{p}_{j}^{t-1}\}$, and over the fine period (short-term horizon) as $\mathcal{H}_{j}^{t}=\{\mathbf{h}_{j}^{1},\mathbf{h}_{j}^{2},...,\mathbf{h}_{j}^{t-1}\}$. We use $t-1$ to constrain access to future interactions and prevent information leakage, i.e., we do not have access to the popularity statistics of $v_{j}$ at time $t$ if we are at time $t$. For example, say an interaction happens on the second Wednesday in February, we consider the coarser and finer time period up until the end of January and the end of the first week in February respectively. To limit computation, we constrain window sizes $m,n$ for $\mathcal{P}$ and $\mathcal{H}$ respectively. Formally, the coarse popularity dynamics of $v_{j}$ at time $t$ is $\mathcal{P}_{j}^{t}=\{\mathbf{p}_{j}^{t-m},\mathbf{p}_{j}^{t-m+1},...,\mathbf{p}_{j}^{t-1}\}$, and the fine popularity dynamics of $v_{j}$ at time $t$ is $\mathcal{H}_{j}^{t}=\{\mathbf{h}_{j}^{t-n},\mathbf{h}_{j}^{t-n+1},...,\mathbf{h}_{j}^{t-1}\}$.
Finally, we compute the embedding of item $v_{j}$ at time $t$ via the universal item representation encoder, defined as a function $\mathcal{E}(\mathcal{P}_{j}^{t},\mathcal{H}_{j}^{t})$ that learns to encode the popularity dynamics $\mathcal{P}_{j}^{t}$ and $\mathcal{H}_{j}^{t}$ into a $d$ dimension vector representation $\mathbf{e}_{j}^{t}$. Specifically, we have:
[…]
4.1.3. Relative Time Interval
We also consider the time interval between two consecutive interactions when modeling sequences. Differences in time intervals might indicate differences in the users’ behaviors. While previous works explore absolute time intervals (Li et al., 2020), different domains exhibit diverse time scales, thus making modeling absolute time intervals ungeneralizable. Therefore, we propose to encode relative time intervals into modeling sequences. Given an interaction sequence $\mathcal{S}_{u}=\{v_{u,1},v_{u,2},...,v_{u,L}\}$ of user $u$, we define the time interval between $v_{u,j}$ and $v_{u,j+1}$ as $t_{u,j}=t(v_{u,j+1})-t(v_{u,j})$, where $t(v_{u,j})$ is the time that user $u$ interacts with item $v_{u,j}$. We then rank the time intervals of user $u$. Define the rank of relative time interval of $t_{u,j}$ as $r_{u,j}=\text{rank}(t_{u,j})$. The relative time interval encoding of interval $t_{u,j}$ is then defined as ${\bm{T}}_{r_{u,j}}\in\mathbb{R}^{d}$, where ${\bm{T}}\in\mathbb{R}^{L\times d}$, following the same setup in (Vaswani et al., 2017), is a fixed sinusoidal encoding matrix defined as:
(3) ${\bm{T}}_{i,2j}=\sin(\frac{i}{L^{2j/d}}),\quad{\bm{T}}_{i,2j+1}=\cos(\frac{i}{L^{2j/d}})$
We also tried a learnable time interval encoding, but it yielded worse performance. We hypothesize that the sinusoidal encoding is more generalizable across domains and the learnable encoding is more prone to overfitting.
4.1.4. Positional Encoding
As we will see in § 4.1.5, the self-attention mechanism does not take the positions of the items into account. Therefore, following (Vaswani et al., 2017), we also inject a fixed positional encoding for each position in a user’s sequence. Denote the positional embedding of a position $l$ as ${\bm{P}}_{l}\in\mathbb{R}^{d}$, where ${\bm{P}}\in\mathbb{R}^{L\times d}$. We compute ${\bm{P}}$ using the same formula in Equation 3. Again, we also tried a learnable positional encoding as presented in (Kang and McAuley, 2018; Sun et al., 2019), but it yielded worse results.
4.1.5. Popularity Dynamics-Aware Transformer
We follow previous works in sequential recommendation (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2020) and propose an extension to the self-attention mechanism by incorporating universal item representations (§ 4.1.2), relative time intervals(§ 4.1.3), and positional encoding (§ 4.1.4).
Firstly, we transform the user sequence $\{v_{u,1},v_{u,2},...,v_{u,|\mathcal{S}_{u}|}\}$ for each user $u$ into a fixed-length sequence $\mathcal{S}_{u}$ = $\{v_{u,1},v_{u,2},...,v_{u,L}\}$ via truncating the oldest interactions or padding, where $L$ is a pre-defined hyper-parameter controlling the maximum length of the sequence. Given a user sequence $\mathcal{S}_{u}=\{v_{u,1},v_{u,2},...,v_{u,L}\}$, we compute its input matrix ${\bm{E}}_{u}$ as:
(4) ${\bm{E}}_{u}=\left[\begin{aligned} &\mathbf{e}^{t}_{u,1}+{\bm{T}}_{r_{u,1}}+{\bm{P}}_{1}\\ &\mathbf{e}^{t^{\prime}}_{u,2}+{\bm{T}}_{r_{u,2}}+{\bm{P}}_{2}\\ &\vdots\\ &\mathbf{e}^{t*}_{u,L}+{\bm{T}}_{r_{u,L}}+{\bm{P}}_{L}\end{aligned}\right]$
$\mathbf{e}^{t}_{u,1},\mathbf{e}^{t^{\prime}}_{u,2},...,\mathbf{e}^{t*}_{u,L}$ is computed from Equation 2, ${\bm{T}}_{r_{u,1}},{\bm{T}}_{r_{u,2}},...,{\bm{T}}_{r_{u,L}}$ and ${\bm{P}}_{1},{\bm{P}}_{2},...,{\bm{P}}_{L}$ are computed following the procedure in § 4.1.3 and § 4.1.4 respectively.
Multi-Head Self-Attention. We adopt a widely used multi-head self-attention mechanism (Vaswani et al., 2017), i.e., Transformers. Specifically, it consists of multiple multi-head self-attention layers (denoted as MHAttn($\cdot$)), and point-wise feed-forward networks (FFN($\cdot$)). The multi-head self-attention mechanism is defined as:
(5) $\begin{split}\mathbf{z}_{u}&=\text{MHAttn}({\bm{E}}_{u})\\ \text{MHAttn}({\bm{E}}_{u})&=\text{Concat}(\text{head}_{1},...,\text{head}_{h}){\bm{W}}^{O}\\ \text{head}_{i}&=\text{Attn}({\bm{E}}_{u}{\bm{W}}_{i}^{Q},{\bm{E}}_{u}{\bm{W}}_{i}^{K},{\bm{E}}_{u}{\bm{W}}_{i}^{V})\end{split}$
where ${\bm{E}}_{u}$ is the input matrix computed from Equation 4, $h$ is a tunable hyper-parameter indicating the number of attention heads, ${\bm{W}}_{i}^{Q},{\bm{W}}_{i}^{K},{\bm{W}}_{i}^{V}\in\mathbb{R}^{d\times d/h}$ are the learnable weight matrices, and $\mathbf{W}^{O}\in\mathbb{R}^{d\times d}$ is also a learnable weight matrix. Attn is the attention function and is formally defined as:
(6) $\begin{split}\text{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})&=\text{softmax}(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d/h}})\mathbf{V}\end{split}$
The scale factor $\sqrt{d/h}$ is used to avoid large values of the inner product, which can lead to numerical instability.
Causality: In sequential recommendation, the prediction of the $L+1$ item should only depend on the first $L$ items that the user has interacted with in the past. However, the $L$-th output of the multi-head self-attention layer contains all the input information. Therefore, as in (Li et al., 2020; Kang and McAuley, 2018), we do not let the model attend to the future items by forbidding links between $Q_{i}$ and $K_{j}$ ($j>i$) in the attention function.
Point-Wise Feed-Forward Network: To add nonlinearity and interactions between different embedding dimensions, we follow previous works in sequential recommendation (Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020) and apply the same point-wise feed-forward network to the output of each multi-head self-attention layer. Formally, suppose the output of the multi-head self-attention layer is $\textbf{z}_{u}$, the point-wise feed-forward network is defined as:
(7) $\begin{split}\text{FFN}(\textbf{z}_{u})&=\text{ReLU}(\textbf{z}_{u}{\bm{W}}_{1}+\mathbf{b}_{1}){\bm{W}}_{2}+\mathbf{b}_{2}\end{split}$
where ${\bm{W}}_{1}\in\mathbb{R}^{d\times d}$ and ${\bm{W}}_{2}\in\mathbb{R}^{d\times d}$ are learnable weight matrices, and $\mathbf{b}_{1}\in\mathbb{R}^{d}$ and $\mathbf{b}_{2}\in\mathbb{R}^{d}$ are learnable bias vectors.
Stacking Layers: As shown in previous works (Kang and McAuley, 2018), stacking multiple multi-head self-attention layers and point-wise feed-forward networks can potentially lead to overfitting and instability during the training. Therefore, we follow previous works (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2020) and apply layer normalization (Ba et al., 2016) and residual connections to each multi-head self-attention layer and point-wise feed-forward network. Formally, we have:
(8) $g(\mathbf{x})=\mathbf{x}+\text{Dropout}(g(\text{LayerNorm}(\mathbf{x})))$
$g(\mathbf{x})$ is either the multi-head self-attention layer or the point-wise feed-forward network. Therefore, for every multi-head self-attention layer and point-wise feed-forward network, we first apply layer normalization to the input, then apply the multi-head self-attention layer or point-wise feed-forward network, and finally apply dropout and add the input $\mathbf{x}$ to the layer output. The LayerNorm function is defined as:
(9) $\text{LayerNorm}(\mathbf{x})=\alpha\odot\frac{\mathbf{x}-\mu}{\sqrt{\sigma^{2}+\epsilon}}+\mathbf{\beta}$
where $\odot$ denotes the element-wise product, $\mu$ and $\sigma$ are the mean and standard deviation of $\mathbf{x}$, $\alpha$ and $\mathbf{\beta}$ are learnable parameters, and $\epsilon$ is a small constant to avoid numerical instability.
4.1.6. Prediction
Given a sequence $\mathcal{S}_{u}$ of user $u$ as input, we denote $\mathbf{q}_{u}$ as the output of the popularity dynamics-aware transformer. Suppose at time $t^{+}$, we want to predict the likelihood of $v_{j}$ being the next item in the sequence, we first compute the item representation $\mathbf{e}_{j}^{t^{+}}$ from § 4.1.2. Then, we predict the score as the inner product of $\mathbf{q}_{u}$ and $\mathbf{e}_{j}^{t^{+}}$, formally:
(10) $\hat{y}(v^{t}_{j}|\mathcal{S}_{u})=<\mathbf{q}_{u},\mathbf{e}_{j}^{t^{+}}>$
Note that there is no information leakage in the prediction process, i.e., we do not assume access to the popularity statistics of $v_{j}$ at time $t^{+}$ if we are at time $t^{+}$ (§ 4.1.2).
4.2. Training Procedure
Now we present how to train the PrepRec model. Similar to (Kang and McAuley, 2018), we adopt the binary cross entropy loss as the objective function, formally:
(11) $\begin{split}\mathcal{L}=-\sum_{\mathcal{S}_{u}\in\mathcal{S}}\sum_{z\in[1,2,...,L-1]}&[\log\sigma(\hat{y}(v^{t}_{z+1}|\mathcal{S}_{u,:z}))\\ &+\sum_{j^{\prime}\notin\mathcal{S}_{u}}\log\sigma(1-\hat{y}(v^{t}_{j^{\prime}}|\mathcal{S}_{u,:z}))]\\ \end{split}$
where $\mathcal{S}_{u,:z}\coloneqq\{v_{u,1},v_{u,2},...,v_{u,z}\}$. $v^{t}_{z+1}$ represents the $z+1$-th item in the sequence that happened at time $t$. We use Adam (Kingma and Ba, 2014) as the optimizer and train the model end-to-end. Note that compared to previous sequential recommenders, we do not have any parameters modeling item IDs. Essentially, we are only optimizing ${\bm{W}}_{p}$ and parameters related to the multi-head self-attention mechanism.
4.3. Zero-shot Inference
Suppose we are given a pre-trained model $\mathcal{F}$ trained on ${\bm{M}}$, where $\mathcal{F}$ is the scoring function learned from source domain ${\bm{M}}$. Denote the interaction matrix of the target domain as ${\bm{M}}^{\prime}$. We first compute the popularity dynamics of each item in ${\bm{M}}^{\prime}$ over a coarser period and a finer period. Then, we apply the pre-trained model $\mathcal{F}$ to ${\bm{M}}^{\prime}$ and compute the prediction score as:
(12) $\hat{y}(v^{t}_{j^{\prime}}|\mathcal{S}^{\prime}_{u})=\mathcal{F}(v^{t}_{j^{\prime}}|\mathcal{S}^{\prime}_{u},{\bm{M}}^{\prime})$
Note that in this procedure, we use the pre-trained model $\mathcal{F}$ trained on domain ${\bm{M}}^{\prime}$ to predict the next item $v^{t}_{j^{\prime}}$ that user $u^{\prime}$ will interact with in domain ${\bm{M}}^{\prime}$. We do not use any auxiliary information in either domain. In addition, none of the parameters in $\mathcal{F}$ are updated during the zero-shot inference process.
To summarize, in this section, we showed how to develop a pre-trained sequential recommender system based on the popularity dynamics of items. We enforce the structure of each interaction in the sequence by the positional encoding and introduce a relative time encoding for modeling time intervals between two consecutive interactions. In addition, we showed the training process and formally defined the zero-shot inference procedure. In the next section, we present experiments to evaluate PrepRec .
5. Experiments
Dataset #users #items #actions avg length density
Office 101,133 27,500 0.74m 7.3 0.03%
Tool 240,464 73,153 1.96m 8.1 0.01%
Movie 70,404 40,210 11.55m 164.2 0.41%
Music 20,539 10,121 0.66m 32.2 0.32%
Epinions 30,989 20,382 0.54m 17.5 0.09%
Table 1. Dataset statistics
We present extensive experiments on five real-world datasets to evaluate the performance of PrepRec, following the problem settings in § 3. We introduce the following research questions (RQ) to guide our experiments: (RQ1) How well can PrepRec perform under the regular, in-domain sequential recommendation setting? (RQ2) Can PrepRec (popularity dynamics) generalize to zero-shot cross-domain and cross-application transfer? (RQ3) How different model components affect the performance of PrepRec ? (RQ4) What is the difference between the predictions of PrepRec and those of the state-of-the-art sequential recommenders?
Dataset Office Tool Movie Music Epinions
SasRec
1.331m
3.581m
2.044m
0.542m
1.054m
BERT4Rec
2.687m
7.233m
4.126m
1.094m
2.127m
TiSasRec
1.367m
3.617m
2.127m
0.578m
1.09m
PrepRec
0.045m
0.045m
0.045m
0.045m
0.045m
Table 2. Comparison of model sizes (i.e., number of learnable parameters in millions) over different datasets. PrepRec uses significantly fewer number of parameters than baselines, ranging from 12 to 90x smaller.
5.1. Datasets and Preprocessing
We evaluate our proposed method on five real-world datasets across different applications, with varying sizes, and density levels.
  • Amazon (Ni et al., 2019) is a series of product ratings datasets obtained from Amazon.com, split by product domain. We consider the Office and Tool product domains in our study.
  • Douban (Song et al., 2019) consists of three datasets across different domains, collected from Douban.com, a Chinese review website. We work with the Movie and Music datasets.
  • Epinions (Tang et al., 2012a, b) is a dataset crawled from product review site Epinions. We utilize the ratings dataset for our study.
We present dataset statistics in Table 1. We compute the density as the ratio of the number of interactions to the number of users times the number of items. Both datasets from Douban (i.e., movie and music) are the densest and have no auxiliary information available, while the Amazon review datasets (i.e., office and tool) are the sparsest.
For fair evaluation, we follow the same preprocessing procedure as previous works (Kang and McAuley, 2018; Sun et al., 2019; Rendle et al., 2010; Tang and Wang, 2018), i.e., we binarize the explicit ratings to implicit feedback. In addition, for each user, we sort interactions by their timestamp and take these chronological sequences as input, using the second most recent action for validation, and the most recent action for testing.
Dataset Office Tool Movie Music Epinions
Metric
R@10
N@10
R@10
N@10
R@10
N@10
R@10
N@10
R@10
N@10
General Recommender Systems
MostPop
0.450
0.272
0.459
0.274
0.586
0.361
0.519
0.327
0.438
0.296
BPR (Rendle et al., 2009)
0.457
0.289
0.363
0.216
0.747
0.477
0.646
0.434
0.568
0.397
NCF (He et al., 2017)
0.446
0.266
0.388
0.239
0.784
0.505
0.652
0.437
0.570
0.396
NGCF (Wang et al., 2019)
0.460
0.291
0.421
0.256
0.788
0.510
0.659
0.441
0.571
0.393
LightGCN (He et al., 2020)
0.465
0.293
0.463
0.275
0.793
0.512
0.665
0.447
0.575
0.396
Sequential Recommender Systems
Caser (Tang and Wang, 2018)
0.512
0.334
0.496
0.297
0.891
0.701
0.796
0.576
0.674
0.475
MARank (Yu et al., 2019)
0.488
0.312
0.468
0.282
0.888
0.683
0.739
0.501
0.628
0.436
SasRec (Kang and McAuley, 2018)
0.539
0.354
0.536
0.337
0.918
0.749
0.816
0.599
$0.705^{*}$
0.501
BERT4Rec (Sun et al., 2019)
$0.541^{*}$
$0.358^{*}$
$0.544^{*}$
$0.350^{*}$
0.900
0.728
$0.816^{*}$
$0.602^{*}$
0.702
$0.512^{*}$
TiSasRec (Li et al., 2020)
0.531
0.349
0.539
0.341
$0.918^{*}$
$0.752^{*}$
0.809
0.523
0.701
0.499
PrepRec
0.536
0.344
0.551
0.339
0.908
0.738
0.782
0.573
0.795
0.580
PrepRec $\Delta$
-0.9%
-3.9%
+1.2%
-3.1%
-1.1%
-1.8%
-1.9%
-4.8%
+12.7%
+13.3%
Interp 0.648 0.483 0.659 0.482 0.929 0.769 0.851 0.653 0.816 0.640
Interp $\Delta$
+19.8%
+34.9%
+21.1%
+35.0%
+1.1%
+2.3%
+4.3%
+8.5%
+15.7%
+25.0%
Table 3. Regular sequential recommendation results. We embolden the best results and mark the best baseline results with $*{}^{\prime}*^{\prime}$. Interp represents the interpolation results between PrepRec and BERT4Rec. PrepRec $\Delta$ denotes the performance difference between PrepRec and the best results among the selected baselines, similar for Interp $\Delta$. Despite not focusing on the performance of regular sequential recommendation, PrepRec achieves comparable performance to the state-of-the-art sequential recommenders, with only on average 0.2% worse than the best performing sequential recommenders in R@10 while having only a fraction of the model size (Table 2). After a simple post-hoc interpolation, we outperform the state-of-the-art sequential recommenders by 13.8% in R@10 on average.
5.2. Baselines
To assess the effectiveness of our proposed method, we compare against the following set of baselines. These baselines include classic general recommendation models and state-of-the-art sequential recommendation models.
  • MostPop is a non-personalized baseline that recommends the most popular items in the training set.
  • BPR (Rendle et al., 2009) learns latent representations for users and items through matrix factorization.
  • NCF (He et al., 2017) learns user and item embeddings through a multi-layer perceptron instead of matrix factorization.
  • NGCF (Wang et al., 2019) uses a bipartite graph to model user-item interactions through a graph convolutional network (Kipf and Welling, 2016).
  • LightGCN (He et al., 2020) is a simplified version of NGCF that removes the feature transformation and nonlinear activation.
  • Caser (Tang and Wang, 2018) learns convolutional filters on user interaction sequences over the sequence order and latent dimensions.
  • MArank (Yu et al., 2019) captures individual and union-level item dependency through multi-order attention.
  • SasRec (Kang and McAuley, 2018) uses self-attention to model user embeddings through its sequence of latent item representations.
  • BERT4Rec (Sun et al., 2019) improves on SasRec by using a bi-directional transformer to capture interactions from both sides.
  • TiSasRec (Li et al., 2020) additionally embeds the raw time interval between consecutive item interactions in the sequence.
To the best of our knowledge, we are the first to tackle the zero-shot sequential recommendation without any auxiliary information. (Hou et al., 2022, 2023; Wang et al., 2023b) are designed for few-shot adaptation and uses auxiliary information. (Ding et al., 2021) is capable of zero-shot recommendation but requires textual descriptions of items, which we do not consider in the experiments. Therefore, we do not include the existing pre-trained sequential recommenders as baselines.
5.3. Experimental Setup
Following previous works (Koren, 2008; He et al., 2017; Kang and McAuley, 2018; Sun et al., 2019), we adopt the leave-one-out evaluation method: for each user, we pair its test item with 100 unobserved items according to the user’s interaction history. Then we rank the test item for the user among the 101 total items. We use two standard evaluation metrics for top-$k$ recommendation: Recall@k (R@k) and Normalized Discounted Cumulative Gain@N (N@k). R@k measures the proportion of test items that appear in the top-$k$ recommendations, and N@k accounts for the ranking position of the test interaction. Our model explicitly utilize popularity information. Therefore, in addition to randomly sampled unobserved items (negatives), we also present results where we sample the negatives based on their popularities, i.e., popular items have higher probabilities of being sampled as negatives, to show that PrepRec can learn discriminative item representations. We report the average of R@k and N@k over all the test interactions.
We use publicly available Pytorch implementations for the baselines. For fair evaluation, we set dimension size $d$ to 50, max sequence length $L$ to $200$, and batch size to $128$ in all models. We use three layers (one base and two additional) in NCF and NGCF, and two attention layers for the sequential models (SasRec, BERT4Rec, TiSasRec, and ours). We use an Adam optimizer and tune the learning rate in the range $\{10^{-4},10^{-3},10^{-2}\}$ and set the weight decay to $10^{-5}$. We use the dropout regularization rate of $0.3$ for all models. We set $\gamma=0.5$ in Equation 1, whose reason we will discuss the reason in  § 5.8.1. We define the coarse and fine period to be 10 and 2 days respectively, and we fix the window size to be $m=12$ and $n=4$ for all datasets (§ 4.1.2). We train PrepRec for a maximum of 80 epochs. All experiments are conducted on a Tesla V100 using PyTorch. We repeat each experiment 5 times with different random seeds and report the average performance.
Dataset Music Office Epinions
Metric
R@10
N@10
R@10
N@10
R@10
N@10
MostPop
0.197
0.139
0.099
0.046
0.163
0.110
SasRec (Kang and McAuley, 2018)
0.749
0.519
0.453
0.291
0.658
0.442
BERT4Rec (Sun et al., 2019)
0.747
0.519
0.461 0.299
0.655
0.456
PrepRec
0.739
0.523
0.443
0.280
0.762 0.551
PrepRec $\Delta$
-1.3%
+0.7%
-2.2%
-6.3%
+15.8%
+17.2%
Table 4. Regular sequential recommendation results with popularity-based negative sampling. We present additional in-domain recommendation results with unobserved (negative) items sampled based on their popularity (Sun et al., 2019) to show that PrepRec can learn discriminative universal item and sequence representations despite depending only on popularity statistics.
[…]
Metric
R@10
N@10
R@10
N@10
R@10
N@10
R@10
N@10
R@10
N@10
Reference: Regular Sequential Recommendation
MostPop
0.450
0.272
0.459
0.274
0.586
0.361
0.519
0.327
0.438
0.296
PrepRec
0.536
0.344
0.551
0.339
0.908
0.738
0.782
0.573
0.795
0.580
[…]
Office $\rightarrow$
0.540 0.326
0.838
0.624
0.755
0.542
0.724
0.512
Tool $\rightarrow$
0.543 0.332
0.881 0.659
0.749
0.536
0.717
0.510
[…]
0.520
0.320
0.508
0.302
0.811 0.600 0.751 0.537
[…]
0.503
0.310
0.496
0.312
0.836
0.636
0.739
0.518
[…]
0.517
0.317
0.470
0.302
0.872
0.656
0.774
0.517
[…]
5.4. Regular Sequential Recommendation (RQ1)
We show the performance of PrepRec against the selected baselines. We want to emphasize that the purpose of this paper is not to outperform the state-of-the-art sequential recommenders but to set a basline for further pre-trained sequential recommenders. We include the results here to investigate how well the popularity dynamics can items and sequences in the sequential recommendation. PrepRec achieves competitive performance- within 2% in R@10 and 5% in N@10, - with the state-of-the-art sequential recommenders. On Epinions, PrepRec even outperforms all the sequential recommenders by 7.3%. This is particularly impressive given that PrepRec has significantly fewer model parameters (Table 2).
Since PrepRec explicitly models popularity information and the MostPop demonstrates competitive performance compared to the remaining baselines, we conduct an additional experiment (Table 4) where we sample the unobserved (negative) items based on their popularity (Sun et al., 2019). As shown in Table 4, MostPop’s performance dropped significantly, while PrepRec shows even more competitive performance on some datasets (e.g., Music and Epinions). This suggests even though PrepRec utilizes non-personalized popularity information, it can still learn discriminative item and sequence representations.
PrepRec learns item representations through popularity dynamics, which is conceptually different from learning representations specific to each item ID. Therefore, we propose a simple post-hoc interpolation to investigate if modeling popularity dynamics can boost the performance of state-of-the-art sequential recommenders. We interpolate the scores from PrepRec with the scores from BERT4Rec as follows: $\hat{y}_{intp}(v^{t}_{j}|\mathcal{S}_{u})=\alpha*\hat{y}_{O}(v^{t}_{j}|\mathcal{S}_{u})+(1-\alpha)*\hat{y}_{S}(v_{j}|\mathcal{S}_{u})$, where $\hat{y}_{O}(v^{t}_{j}|\mathcal{S}_{u})$ and $\hat{y}_{S}(v_{j}|\mathcal{S}_{u})$ are the scores from PrepRec (Equation 10) and BERT4Rec, respectively. We set $\alpha=0.5$ for all datasets. After interpolation, there is a significant performance boost by up to 34.9% in N@10. This suggests that our model effectively captures dynamic item properties that are missing with the static item embeddings of SOTA sequential models. Gains are largest in the medium and low-density datasets (Epinions, Amazon), indicating that our model is particularly complimentary to existing methods in sparse datasets where item embeddings are less informative.
5.5. Zero-shot Transfer(RQ2)
We follow the zero-shot inference setting introduced in § 4.3 and report the results in Table 5. We also include the results of the regular sequential recommendation for reference. In the zero-shot setting, after training on source datasets, PrepRec shows minimal performance reduction in the target datasets (i.e., 6% maximum and 2% average reduction in R@10). The best zero-shot transfer results from PrepRec only fall short against the selected sequential recommendation baselines (Table 3) by up to 4% and even outperform (by up to 6.5%) them on the target datasets (e.g., Epinions and Office). We found that our models trained on Movie and tools show the highest generalizability, even outperforming the target-trained models on Music (0.811 vs. 0.782 on R@10). We conjecture that this is because Movie is the largest dataset in terms of the number of interactions. Overall, these results show PrepRec ’s effectiveness in zero-shot transfer without any training on interaction data or side information. In addition, this experiment also demonstrates that the popularity dynamics-based item and sequence representations are generalizable across domains and applications.
5.6. Ablation Study (RQ3)
Dataset Music Office Epinions
Metric
R@10
N@10
R@10
N@10
R@10
N@10
PrepRec 0.782 0.573
0.536
0.344 0.795 0.580
w/o Relative Time ${\bm{T}}$ (§ 4.1.3)
0.734
0.514
0.541
0.334
0.782
0.562
w/o Positional ${\bm{P}}$ (§ 4.1.4)
0.765
0.544
0.530
0.332
0.772
0.554
Sinuoisal Popularity Encoding
0.779
0.570
0.529
0.340
0.772
0.561
Table 6. Ablation study of PrepRec ’s different variants.
Here, we assess the importance of different components crucial to PrepRec , i.e., relative time encoding (§ 4.1.3), positional encoding (§ 4.1.4), and popularity encoder $E_{p}$ (§ 4.1.1). We find that removing relative time encoding ${\bm{T}}$ results in the largest performance drop on both the Music and Office datasets. This suggests that the relative time encoding is crucial for effectively capturing the popularity dynamics. Removing positional encoding ${\bm{P}}$ results in a maximum of 2.2% drop in R@10 on the Office dataset, indicating positional encoding is important for capturing sequential information. Finally, changing $E_{p}$ to the non-linear sunusoidal encoding shows worse performance on all datasets, meaning that the linear encoding is more suitable for capturing the popularity dynamics.
5.7. Qualitative Analysis(RQ4)
Figure 2. Recommendation results for different item popularity groups. Item Popularity Group 1 represents the least popular items, and Group 5 represents the most popular items. PrepRec achieves better performance on long-tail items, while the sequential recommenders show stronger performance on popular items.
Here we present a detailed analysis regarding the performance of PrepRec . We separate test items into equally sized groups based on their popularity in the training set. We then compute the average R@10 and N@10 for each group. We present the results in Figure 2. We find that PrepRec achieves better performance on item group with the least interactions, i.e., long-tail items, while the SasRec and BERT4Rec show stronger performance on popular items. Long-tail item recommendation is a particularly challenging task explored by many previous works (Sankar et al., 2021) and requires recommenders able to learn high-quality representations with just a few interactions. This corresponds to our observation that PrepRec is more robust to data sparsity and can learn discriminative item and sequence representations (§ 5.4), showing that long-tail item recommendation can benefit from PrepRec ’s popularity dynamics-based item representations.
5.8. Further Analysis
Dataset Music Office Epinions
Metric
R@10
N@10
R@10
N@10
R@10
N@10
$\gamma=0.5^{*}$ (weighted -pop)
0.782 0.573 0.536 0.344 0.795 0.580
$\gamma=0$ (Curr-pop)
0.749
0.542
0.512
0.328
0.689
0.496
$\gamma=1$ (cumul-pop)
0.695
0.452
0.530
0.330
0.733
0.505
Table 7. Recommendation results for varying the discounting factor $\gamma$ in § 4.1.2. $\gamma=0.5$ is the default setting, denoted by $*{}^{\prime}*^{\prime}$. We find that $\gamma=0.5$ generally outperforms the other two settings
5.8.1. Effect of discounting factor $\gamma$
We examine the effect of different preprocessing weights $\gamma$ used in popularity calculation (§ 4.1.1). In particular, $\gamma=1$ corresponds to the cumulative popularity, or in other words, at a given time period $t$, the overall number of interactions up to period $t$. On the other hand, $\gamma=0$ corresponds to the current popularity, or percentiles are calculated over interactions just in $t$, same as $b_{j}^{t}$ in Equation 1. When $\gamma=0.5$, it can be interpreted as interactions being exponentially weighted by time, with a half-life of 1 time period. We find that $\gamma=0.5$ outperforms the other two settings, with the largest gains of around 12% R@10 and 27% N@10 over cumul-pop in the dense Music dataset, and the largest gains over curr-pop in the sparser Office (5% N@10 and 4% N@10) and Epinions ($15\%$ R@10 and 17%N@10) datasets. We suspect this is due to cumulative measures in denser datasets failing to capture recent trends due to the large historical presence, while current-only measures in sparser datasets convey too little or noisy information and lose the information of long-term trends. curr-pop shows decent performance on the Music dataset, suggesting that Music trends might be more cyclical and thus the current popularity is more informative.
[…]
Metric
R@10
N@10
R@10
N@10
R@10
N@10
PrepRec
0.803
0.591
0.472 0.300 0.489 0.264
SasRec
0.815
0.599
0.437
0.290
0.433
0.245
BERT4Rec
0.816 0.602
0.407
0.249
0.433
0.255
Table 8. Recommendation results for fine-tuning PrepRec. We fine-tune PrepRec and retrain the baselines from scratch on the target dataset.
5.8.2. Fine-tune Capability
While we focus on zero-shot cross-domain and cross-application transfers, we also investigate PrepRec’s fine-tune capability. To ensure target datasets are smaller than the source, we further process the target datasets such that they are no more than 10% of the source datasets’ total interactions. After further processing, we follow the same experimental setup in § 5.3. We fine-tune PrepRec and retrain the baselines from scratch on the target dataset and report the results in Table 8. We find that PrepRec , after fine-tuning, outperforms the selected baselines on Office and Epinions by up to 12.9%. This indicates that PrepRec is capable of learning from the limited data and can be further fine-tuned to achieve better performance, which further expands the use case of PrepRec .
Dataset Music Office Epinions
Metric
R@10
N@10
R@10
N@10
R@10
N@10
Fine:2 days; Coarse:10 days $*{}^{*}$
0.782 0.573
0.536
0.344 0.795 0.580
Fine:4 days; Coarse:15 days
0.778
0.553
0.537
0.341
0.790
0.574
Fine:7 days; Coarse:30 days
0.760
0.509
0.526
0.334
0.757
0.543
Table 9. Recommendation results for varying time horizons. Fine and coarse time horizons are used for short-term and long-term popularity dynamics respectively (§ 4.1.1).
5.8.3. Effect of Different Time Horizons
We study the effect of different time horizons to PrepRec . We found that in general, long-term horizons of 30 days and short-term horizons of 7 days perform worse than the other settings. This is likely because the long-term horizon might lead to the lack of resolutions in popularity statistics. We also find that depending on the dataset, the effect of different time horizons also varies. For example, both Music and Epinions show larger performance decrease from short to long-term horizons than Office. This could be because Music and Epinions are more sensitive to recent trends than Office, or their data are denser in terms of time granularity.
Dataset Music Office Epinions
Metric
R@10
N@10
R@10
N@10
R@10
N@10
0.782 0.573 0.536 0.344 0.795 0.580
t-2
0.775
0.553
0.530
0.338
0.728
0.516
t-6
0.730
0.497
0.523
0.326
0.682
0.473
t-12
0.705
0.462
0.507
0.313
0.645
0.445
Table 10. Recommendation results for varying prediction time. Where $t-n$ denotes $n$ finer time periods ago. $t-1$ is the default setting, denoted by $*{}^{\prime}*^{\prime}$. The performance drops as $n$ goes larger, indicating the importance of recent trends in sequential recommendation.
5.8.4. Effect of Prediction Time
In PrepRec, we use the popularity trends up to some time as the context to construct universal item representations. To prevent information leakage, we can only use the popularity dynamics computed prior to the time of the interaction, as illustrated in § 4.1.2 and § 4.1.6. Intuitively, we should compute popularity dynamics as close to the actual interaction time as possible. However, experimenting with this can reveal the importance of recent trends in model performance ( Table 10). Thus, denoting $t-n$ to be $n$ fine time periods ago. We round the long-term horizon according to the change in $t$. Per expectation, we find that $t-6$ and $t-12$ drop performance significantly across all datasets while $t-2$ performs relatively close to $t-1$, revealing that for the experimented datasets, capturing the latest popularity changes is crucial.
5.9. Discussion
PrepRec shows excellent potential as a pre-trained sequential recommendation model, with competitive performance to the best baseline with a fraction of the model size, and is capable of zero-shot transfer across domains and applications. We argue that PrepRec is particularly useful in the following scenarios: 1) initial sequential model when the data in the domain is sparse; 2) backbone for developing more complex sequential recommenders (i.e., prediction interpolation) 3) online recommendation settings. PrepRec does not explicitly model Item IDs, it can also easily generalize to new items/users with no retraining by just updating the popularity dynamics. We leave this as future work.
6. Conclusion
Inspired by the impact of pre-trained models, we propose PrepRec , a pre-trained sequential recommendation model that is capable of cross-domain and cross-application zero-shot transfer. We assume no access to auxiliary information (e.g., textual data), and we set a baseline for pre-trained sequential recommenders under the most challenging setting. PrepRec learns to represent items based on their popularity dynamics and computes sequence representations via a popularity dynamics-aware transformer. Through extensive experiments, we empirically show that PrepRec can achieve comparable performance to or even outperform state-of-the-art sequential recommenders on regular sequential recommendation tasks. PrepRec also shows excellent generalizability and can zero-shot transfer to a new domain with minimal performance reduction, indicating that popularity dynamics play an important role in sequential recommendation and are crucial for developing generalizable sequential recommenders.