### v3

# A Pre-trained Zero-shot Sequential Recommendation Framework via Popularity Dynamics

Abstract.
This paper proposes a novel pre-trained framework for zero-shot cross-domain sequential recommendation without auxiliary information. While using auxiliary information (e.g., item descriptions) seems promising for cross-domain transfer, a cross-domain adaptation of sequential recommenders can be challenging when the target domain differs from the source domain—item descriptions are in different languages; metadata modalities (e.g., audio, image, and text) differ across source and target domains. If we can learn universal item representations independent of the domain type (e.g., groceries, movies), we can achieve zero-shot cross-domain transfer without auxiliary information. Our critical insight is that user interaction sequences highlight shifting user preferences via the popularity dynamics of interacted items. We present a pre-trained sequential recommendation framework: PrepRec, which utilizes a novel popularity dynamics-aware transformer architecture. Through extensive experiments on five real-world datasets, we show that PrepRec, without any auxiliary information, can zero-shot adapt to new application domains and achieve competitive performance compared to state-of-the-art sequential recommender models. In addition, we show that PrepRec complements existing sequential recommenders. With a simple post-hoc interpolation, PrepRec improves the performance of existing sequential recommenders on average by 11.8% in Recall@10 and 22% in NDCG@10. We provide an anonymized implementation of PrepRec at https://github.com/CrowdDynamicsLab/preprec.
Recommender System, Zero-shot Sequential Recommendation
journalyear: 2024copyright: rightsretainedconference: 18th ACM Conference on Recommender Systems; October 14–18, 2024; Bari, Italybooktitle: 18th ACM Conference on Recommender Systems (RecSys ’24), October 14–18, 2024, Bari, Italydoi: 10.1145/3640457.3688145isbn: 979-8-4007-0505-2/24/10ccs: Information systems Recommender systems
1. INTRODUCTION
Modeling sequential user behavior is critical to the success of online applications such as e-commerce, video streaming, and social media. Despite essential innovations for tackling the sequential recommendation task (Rendle et al., 2010; Hidasi et al., 2015; Tan et al., 2016; Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020), these systems have some limitations. Firstly, they must be trained from scratch for each application domain because they learn domain-specific item representations (Sun et al., 2019; Kang and McAuley, 2018), which is resource-consuming and limits model reuse across domains. Even within a domain, they must be retrained when there is a large influx of users or items to maintain performance. Prior work tackles these limitations by incorporating auxiliary information (Ding et al., 2021; Hou et al., 2022; Hao et al., 2021), e.g., item descriptions. However, using auxiliary information can be problematic for cross-domain transfer if item descriptions are in different languages (e.g., English and Chinese), or if the metadata modalities (e.g., , audio, image, and text) differ across domains.
This paper tackles a challenging cross-domain transfer setting where we assume no access to auxiliary information. Thus, we ask: can we build a pre-trained sequential recommender system capable of cross-domain and cross-application transfer without ny auxiliary information? (e.g., using the model trained for online shopping in the US to predict the next movie a user in China will watch). Our work is a performance baseline for cross-domain tasks; using compatible (i.e., same language/modality) auxiliary information across domains, can only improve the performance.
Figure 1. Jensen-Shannon divergence between two consecutive windows ($k$, $k+1$), where we measure the change in popularity of items in the user’s sequence. The sampled users are from the Amazon Office dataset. This shows that there exist temporal item popularity shifts in user interaction sequences. We set the window size to 10 and stride to 5.
At first glance, developing pre-trained sequential recommenders for cross-domain inference seems impossible. While we see pre-trained language (Devlin et al., 2018; Brown et al., 2020; OpenAI, 2023; Raffel et al., 2020; Liu et al., 2019) and vision models (Dosovitskiy et al., 2020; He et al., 2016; Radford et al., 2021) show excellent generalizability across datasets and applications, being able to achieve state-of-the-art performance by just a few fine-tuning steps (Devlin et al., 2018; Liu et al., 2019) or even without any training (OpenAI, 2023; Brown et al., 2020) (i.e., zero-shot transfer), there are essential differences. The representations learned by the pre-trained language model seem universal since the training domain and the application domain (e.g., text prediction and generation) share the same language and vocabulary, supporting the effective reuse of the word representations. However, in the cross-domain recommendation, the items are distinct across domains in recommendation datasets (e.g., grocery items vs movies). Therefore, forming such generalizable correspondence is nearly impossible if we learn representations for each item within each domain. Recent work explores pre-trained models for sequential recommendation (Ding et al., 2021; Hou et al., 2022; Hao et al., 2021) within the same application (e.g., online retail). However, they assume access to metadata of items (e.g., item description), which is domain-dependent and is often not generalizable to other domains. These models cannot learn universal representations of items; instead, they bypass the representation learning problem by using additional item-side information.
Our Insight: There exist item popularity shifts in the user’s sequence, as indicated in Figure 1. The item popularity shifts can be explained as temporal shifts in the user’s preferences. For example, a user might be interested in buying some common office goods such as pens, papers, and notebooks, but afterward, they might look for other less common office goods such as a whiteboard or a desk. Previous works try to learn users’ preference from the past sequence but ignore the crucial aspect of item popularity dynamics, which could indicate the user’s changing preferences. We know that the marginal distribution of user and item activities are heavy-tailed across datasets, supported by prior work in network science (Barabási and Albert, 1999; Barabasi, 2005) and by experiments in recommender systems (Salganik et al., 2006). In addition, recent work in recommender systems suggests that the popularity dynamics of items are also crucial for predicting users’ behaviors (Ji et al., 2020).
Present Work: In this paper, we develop universal, transferable item representations for the zero-shot, cross-domain setting based on the popularity dynamics of items. We explicitly model the popularity dynamics of items and propose a novel pre-trained sequential recommendation framework: PrepRec. We learn universal item representations based on their popularity dynamics instead of their explicit item IDs or auxiliary information. We encode the relative time interval between two consecutive interactions via relative-time encoding and the position of each interaction in the sequence by positional encoding. Using physical time ensures that the predictions are not anti-causal, i.e., using the future interactions to predict the present. We propose a popularity dynamics-aware transformer architecture for learning universal sequence representations. We show that it is possible to build a pre-trained sequential recommender system capable of cross-domain and cross-application transfer without any auxiliary information. Our key contributions are as follows:
Universal item and sequence representations::
We are the first to learn universal item and sequence representations for sequential recommendation without any auxiliary information by exploiting item popularity dynamics. In contrast, prior research learns item representations for each item ID or through item auxiliary information. We learn universal item representations by modeling item popularity dynamics of two temporal resolutions: coarse and fine-grained. We learn universal sequence representations using a carefully designed popularity dynamics-aware transformer architecture. These universal item and sequence representations make possible pre-trained sequential recommender systems capable of cross-domain and cross-application transfer without any auxiliary information.
Zero-shot transfer without auxiliary information::
We propose a new challenging setting for pre-trained sequential recommender systems: zero-shot cross-domain and cross-application transfer without any auxiliary information. In contrast, previous pre-trained sequential recommenders requires overlapping users (Yuan et al., 2020), application-dependent auxiliary information (Ding et al., 2021; Hou et al., 2022; Hao et al., 2021; Hou et al., 2023; Wang et al., 2023b), and are few-shot adapted to related domains within the same application (Hou et al., 2022, 2023; Wang et al., 2023b). Our work establishes a performance baseline for cross-domain sequential recommenders that use compatible (i.e., same language/modality) auxiliary information across domains, as such metadata can only improve the performance of cross-domain transfer.
With extensive experiments, we empirically show that PrepRec has excellent generalizability across domains and applications. Remarkably, had we trained a state-of-the-art model from scratch for the target domain, instead of zero-shot transfer using PrepRec, the maximum performance gain over PrepRec would have been only 4%. In addition, we show that PrepRec is complementary to state-of-the-art sequential recommenders and with a post-hoc interpolation, PrepRec can outperform the state-of-the-art sequential recommender system on average by 11.8% in Recall@10 and 22% in NDCG@10. We attribute the improvements to the performance gains over long-tail items, which we show in the qualitative analysis. With this work, we set a baseline for pre-trained sequential recommenders and show that popularity dynamics not only enable us to build a pre-trained sequential recommender system capable of zero-shot transfer but also significantly boost the performance of sequential recommendation.
2. RELATED WORK
Sequential Recommendation: Sequential recommenders model user behavior as a sequence of interactions, and aim to predict the next item that a user will interact with. Early sequential recommenders adopt Markov chains (Rendle et al., 2010; Shani et al., 2005) and basic neural network architectures (Tang and Wang, 2018; Tuan and Phuong, 2017; Hidasi et al., 2016, 2015; Tan et al., 2016). With the success of Transformer (Vaswani et al., 2017) in modeling sequential data  (Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020)adopt the transformer architecture for sequential recommendation. Additionally, (Li et al., 2020) considers the timestamps of each interaction and proposes a time-aware attention mechanism.  (Lv et al., 2019; Ying et al., 2018; Tan et al., 2021) separate interaction sequences and categorize them to show the long-term and short-term interests of users. Temporal sequential recommenders (Koren, 2009; Wu et al., 2017; Zhang et al., 2014) models the change in users’ preferences. These works, while achieving state-of-the-art performance, only focus on the regular sequential recommendation and cannot transfer to other domains.
Cross-domain Recommendation: Cross-domain recommendation literature leverages the information-rich domain to improve the recommendation performance on the data-sparse domain (Hu et al., 2018; Li and Tuzhilin, 2020; Man et al., 2017). However, most of these works assume user or item overlap (Zhu et al., 2021; Zhao et al., 2020; Hu et al., 2018; Li and Tuzhilin, 2020; Man et al., 2017) for effective knowledge transfer. Other cross-domain literature focuses on the cold-start problem (Lu et al., 2020; Zhu et al., 2021; Feng et al., 2021; Volkovs et al., 2017; Wei et al., 2021; Du et al., 2020; Liu et al., 2021; Lee et al., 2019; Dong et al., 2020). In addition, multi-domain recommenders  (Sheng et al., 2021; Ariza-Casabona et al., 2023) leverage multi-domain data to gain insights into user preferences and item characteristics.
Pre-trained Sequential Recommenders: Recently, pre-trained recommenders have caught the attention of the community. ZESRec (Ding et al., 2021) is capable of zero-shot sequential recommendations. However, it only works for closely related domains and requires item metadata. PeterRec (Yuan et al., 2020) requires overlapping users in both domains. On the other hand, finetuning-based models, e.g., MISSRec (Wang et al., 2023b), UnisRec (Hou et al., 2022), and VQ-Rec (Hou et al., 2023), are not designed for zero-shot sequential recommendation and works within the same application (e-commerce), and they rely on application-dependent auxiliary information.  (Wang et al., 2023a) investigates the joint and marginal activity distribution of users and items, but are not suitable for the sequential recommendation task.
To summarize, prior works on sequential recommendation focus on learning high-quality representations for each item in the training set and are not generalizable across domains. Pre-trained sequential recommenders are evaluated on closely related domains and platforms and rely heavily on application-dependent auxiliary information of items.
3. PROBLEM DEFINITION
In this section, we formally define the research problems this paper addresses (i.e., regular sequential recommendation and zero-shot sequential recommendation) and introduce our notations.
In sequential recommendation, denote ${\bm{M}}$ as the implicit feedback matrix, $\mathcal{U}$ = $\{u_{1},u_{2},...,u_{|\mathcal{U}|}\}$ as the set of users, $\mathcal{V}$ = $\{v_{1},v_{2},...,v_{|\mathcal{V}|}\}$ as the set of items. The goal of sequential recommendation is to learn a scoring function, that predicts the next item $v_{u,t}$ given a user $u$ ’s interaction history $\mathcal{S}_{u}=\{v_{u,1},v_{u,2},...,v_{u,t-1}\}$. Note that in this paper, since we model time explicitly, we assume access to the timestamp of each interaction, including the next item interaction. We argue that this is a reasonable assumption since the timestamp of the next interaction is always available in practice. For example, if Alice logs in to Netflix, Netflix will always know when Alice logs in and can predict the next movie for Alice. Formally, we define the scoring function as $\mathcal{F}(v^{t}|\mathcal{S}_{u},{\bm{M}})$, where $t$ is the time of the prediction.
Zero-shot Sequential Recommendation: Given two domains ${\bm{M}}$ and ${\bm{M}}^{\prime}$ over $\mathcal{U},\mathcal{V}$ and $\mathcal{U}^{\prime},\mathcal{V^{\prime}}$ respectively, we study the zero-shot recommendation problem in the scenario where the domains are different (${\bm{M}}\cap{\bm{M}}^{\prime}=\varnothing$), users are disjoint ($\mathcal{U}\cap\mathcal{U^{\prime}}=\varnothing$), and item sets are unique ($\mathcal{V}\cap\mathcal{V^{\prime}}=\varnothing$). The goal is to produce a scoring function $\mathcal{F}^{\prime}$ without training on ${\bm{M}}^{\prime}$ directly. In other words, the scoring function $\mathcal{F}^{\prime}$ has to be trained on a different interaction matrix ${\bm{M}}$. Furthermore, we assume there is no metadata associated with users or items, which makes the problem particularly challenging but crucial to study. We want to set a baseline for pre-trained sequential recommenders, using metadata can only improve and simply the problem.
4. PREPREC FRAMEWORK
Figure 2. Model Architecture of PrepRec
We first introduce the model architecture of PrepRec (§ 4.1) then the training procedure (§ 4.2). Finally, we formally define the zero-shot inference process (§ 4.3).
4.1. Model Architecture
The first step of building a pre-trained sequential recommender is to learn universal item representations. Our solution is to exploit the item popularity statistics to learn universal item representations. We learn to represent items at a given timestamp through the changes in their popularity histories over different periods, i.e., popularity dynamics. We propose a popularity dynamics-aware Transformer architecture that obtains the representation of users’ behavior sequences through item popularity dynamics.
4.1.1. Item Popularity Encoder
We learn to represent items based on their popularity dynamics, i.e., changes in their popularity histories. Intuitively, popularity can be calculated by two horizons: long-term and short-term. Long-term horizons reflect the overall popularity of items, whereas short-term horizons should capture the recent trends in the domain. For example, the long-term popularity of a winter coat measures how popular is the coat in general, while its short-term popularity reflects more temporal changes, e.g., season, weather conditions, and fashion trends. Therefore, consider an item $v_{j}$ that has interaction at time $t$, denoted as $v_{j}^{t}$, we define two popularity representations for $v_{j}^{t}$: popularity $\mathbf{p}_{j}^{t}\in\mathbb{R}^{k}$ over a coarse period (e.g., month) and popularity $\mathbf{h}_{j}^{t}\in\mathbb{R}^{k}$ over a fine period (e.g., week).
To calculate $\mathbf{p}_{j}^{t}$ and $\mathbf{h}_{j}^{t}$, we first calculate the popularities of $v_{j}^{t}$ over the two horizons, denoted as $a_{j}^{t}\in\mathbb{R}^{+}$ (coarse period number of interactions) and $b_{j}^{t}\in\mathbb{R}^{+}$ (fine period number of interactions). Specifically, we calculate them as:
(1) $a_{j}^{t}=\sum_{m=1}^{t}\gamma^{t-m}c_{a}(v_{j}^{m}),\quad b_{j}^{t}=c_{b}(v_{j}^{t})$
where $\gamma\in\mathbb{R}^{+}$ is a pre-defined discount factor and $c_{a}(v_{j}^{m})$ is the number of interactions of $v_{j}$ over a coarse time period $m$. Similarly, $c_{b}(v_{j}^{t})$ denotes the number of interactions of $v_{j}$ over a fine period $t$. We do not impose the discounting factor when computing $b_{j}^{t}$ since we want it to capture the current popularity information, whereas $a_{j}^{t}$ captures the cumulative popularity of an item over a longer horizon.
To make item popularity comparable across domains, we calculate the percentiles of $a_{j}^{t}$ and $b_{j}^{t}$ relative to their corresponding coarser and finer popularity distributions over all items at time $t$, denoted as $P(a_{j}^{t})\in\mathbb{R}^{+}$ and $P({b_{j}^{t}})\in\mathbb{R}^{+}$, respectively.
We now encode the popularity percentiles $P(a_{j}^{t})$ and $P({b_{j}^{t}})$ into $k$ dimensional vector representations $\mathbf{p}_{j}^{t}$ and $\mathbf{h}_{j}^{t}$ respectively. Denote the popularity encoder as $E_{p}:\mathbb{R}^{+}\rightarrow\mathbb{R}^{k}$, which takes in a percentile value. Suppose given the popularity percentile $P(a_{j}^{t})\in\mathbb{R}^{+}$ over a coarse time period $t$, the coarse level popularity vector representation $\mathbf{p}_{j}^{t}\in\mathbb{R}^{k}$ is computed as follows:
$\mathbf{p}_{j}^{t}$ $=E_{p}(P(a_{j}^{t}))$
$(\mathbf{p}_{j}^{t})_{i}$ =$=\begin{cases}1-\{\frac{P}{k-1}\},&\text{if }i=\lfloor\frac{P}{k-1}\rfloor\\ \{\frac{P}{k-1}\},&\text{if }i=\lfloor\frac{P}{k-1}\rfloor+1\\ 0,&\text{otherwise}\end{cases}$
[…]
4.1.2. Universal Item Representation
We now define the popularity dynamics of $v_{j}$ at time $t$ over the coarse period (long-term horizon) to be $\mathcal{P}_{j}^{t}=\{\mathbf{p}_{j}^{1},\mathbf{p}_{j}^{2},...,\mathbf{p}_{j}^{t-1}\}$, and over the fine period (short-term horizon) as $\mathcal{H}_{j}^{t}=\{\mathbf{h}_{j}^{1},\mathbf{h}_{j}^{2},...,\mathbf{h}_{j}^{t-1}\}$. We use $t-1$ to constrain access to future interactions and prevent information leakage, i.e., we do not have access to the popularity statistics of $v_{j}$ at time $t$ if we are at time $t$. For example, say an interaction happens on the second Wednesday in February, we consider the coarser and finer time period up until the end of January and the end of the first week in February respectively. To limit computation, we constrain window sizes $m,n$ for $\mathcal{P}$ and $\mathcal{H}$ respectively. Formally, the coarse popularity dynamics of $v_{j}$ at time $t$ is $\mathcal{P}_{j}^{t}=\{\mathbf{p}_{j}^{t-m},\mathbf{p}_{j}^{t-m+1},...,\mathbf{p}_{j}^{t-1}\}$, and the fine popularity dynamics of $v_{j}$ at time $t$ is $\mathcal{H}_{j}^{t}=\{\mathbf{h}_{j}^{t-n},\mathbf{h}_{j}^{t-n+1},...,\mathbf{h}_{j}^{t-1}\}$.
Finally, we compute the embedding of item $v_{j}$ at time $t$ via the universal item representation encoder, defined as a function $\mathcal{E}(\mathcal{P}_{j}^{t},\mathcal{H}_{j}^{t})$ that learns to encode the popularity dynamics $\mathcal{P}_{j}^{t}$ and $\mathcal{H}_{j}^{t}$ into a $d$ dimension vector representation $\mathbf{e}_{j}^{t}$. Specifically, we have:
(2) $\mathbf{e}_{j}^{t}=\mathcal{E}(\mathcal{P}_{j}^{t},\mathcal{H}_{j}^{t})={\bm{W}}_{p}[(\|_{i=t-m}^{t-1}\mathbf{p}_{j}^{i})\|(\|_{i=t-n}^{t-1}\mathbf{h}_{j}^{i})]$
[…]
In addition, we define $\|_{i=t-m}^{t-1}\mathbf{p}_{j}^{i}\coloneq\mathbf{p}_{j}^{t-m}\|\mathbf{p}_{j}^{t-m+1}\|...\|\mathbf{p}_{j}^{t-1}$ and $\|_{i=t-n}^{t-1}\mathbf{h}_{j}^{i}\coloneq\mathbf{h}_{j}^{t-n}\|\mathbf{h}_{j}^{t-n+1}\|...\|\mathbf{h}_{j}^{t-1}$. The item popularity dynamics encoder can effectively capture the popularity change of items over different time periods. Most importantly, it does not take explicit item IDs or auxiliary information as input to compute the item embeddings. Instead, it learns to represent items through their popularity dynamics, which is universal across domains and applications.
4.1.3. Relative Time Interval
We also consider the time interval between two consecutive interactions when modeling sequences. Differences in time intervals might indicate differences in the users’ behaviors. While previous works explore absolute time intervals (Li et al., 2020), different domains exhibit diverse time scales, thus making modeling absolute time intervals ungeneralizable. Therefore, we propose to encode relative time intervals into modeling sequences. Given an interaction sequence $\mathcal{S}_{u}=\{v_{u,1},v_{u,2},...,v_{u,L}\}$ of user $u$, we define the time interval between $v_{u,j}$ and $v_{u,j+1}$ as $t_{u,j}=t(v_{u,j+1})-t(v_{u,j})$, where $t(v_{u,j})$ is the time that user $u$ interacts with item $v_{u,j}$. We then rank the time intervals of user $u$. Define the rank of relative time interval of $t_{u,j}$ as $r_{u,j}=\text{rank}(t_{u,j})$. The relative time interval encoding of interval $t_{u,j}$ is then defined as ${\bm{T}}_{r_{u,j}}\in\mathbb{R}^{d}$, where ${\bm{T}}\in\mathbb{R}^{L\times d}$, following the same setup in (Vaswani et al., 2017), is a fixed sinusoidal encoding matrix defined as:
(3) ${\bm{T}}_{i,2j}=\sin(\frac{i}{L^{2j/d}}),\quad{\bm{T}}_{i,2j+1}=\cos(\frac{i}{L^{2j/d}})$
We also tried a learnable time interval encoding, but it yielded worse performance. We hypothesize that the sinusoidal encoding is more generalizable across domains and the learnable encoding is more prone to overfitting.
4.1.4. Positional Encoding
As we will see in § 4.1.5, the self-attention mechanism does not take the positions of the items into account. Therefore, following (Vaswani et al., 2017), we also inject a fixed positional encoding for each position in a user’s sequence. Denote the positional embedding of a position $l$ as ${\bm{P}}_{l}\in\mathbb{R}^{d}$, where ${\bm{P}}\in\mathbb{R}^{L\times d}$. We compute ${\bm{P}}$ using the same formula in Equation 3. Again, we also tried a learnable positional encoding as presented in (Kang and McAuley, 2018; Sun et al., 2019), but it yielded worse results.
4.1.5. Popularity Dynamics-Aware Transformer
We follow previous works in sequential recommendation (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2020) and propose an extension to the self-attention mechanism by incorporating universal item representations (§ 4.1.2), relative time intervals(§ 4.1.3), and positional encoding (§ 4.1.4).
Firstly, we transform the user sequence $\{v_{u,1},v_{u,2},...,v_{u,|\mathcal{S}_{u}|}\}$ for each user $u$ into a fixed-length sequence $\mathcal{S}_{u}$ = $\{v_{u,1},v_{u,2},...,v_{u,L}\}$ via truncating the oldest interactions or padding, where $L$ is a pre-defined hyper-parameter controlling the maximum length of the sequence. Given a user sequence $\mathcal{S}_{u}=\{v_{u,1},v_{u,2},...,v_{u,L}\}$, we compute its input matrix ${\bm{E}}_{u}$ as:
(4) ${\bm{E}}_{u}=\left[\begin{aligned} &\mathbf{e}^{t}_{u,1}+{\bm{T}}_{r_{u,1}}+{\bm{P}}_{1}\\ &\mathbf{e}^{t^{\prime}}_{u,2}+{\bm{T}}_{r_{u,2}}+{\bm{P}}_{2}\\ &\vdots\\ &\mathbf{e}^{t*}_{u,L}+{\bm{T}}_{r_{u,L}}+{\bm{P}}_{L}\end{aligned}\right]$
$\mathbf{e}^{t}_{u,1},\mathbf{e}^{t^{\prime}}_{u,2},...,\mathbf{e}^{t*}_{u,L}$ is computed from Equation 2, ${\bm{T}}_{r_{u,1}},{\bm{T}}_{r_{u,2}},...,{\bm{T}}_{r_{u,L}}$ and ${\bm{P}}_{1},{\bm{P}}_{2},...,{\bm{P}}_{L}$ are computed following the procedure in § 4.1.3 and § 4.1.4 respectively.
Multi-Head Self-Attention. We adopt a widely used multi-head self-attention mechanism (Vaswani et al., 2017), i.e., Transformers. Specifically, it consists of multiple multi-head self-attention layers (denoted as MHAttn($\cdot$)), and point-wise feed-forward networks (FFN($\cdot$)). The multi-head self-attention mechanism is defined as:
(5) $\begin{split}\mathbf{z}_{u}&=\text{MHAttn}({\bm{E}}_{u})\\ \text{MHAttn}({\bm{E}}_{u})&=\text{Concat}(\text{head}_{1},...,\text{head}_{h}){\bm{W}}^{O}\\ \text{head}_{i}&=\text{Attn}({\bm{E}}_{u}{\bm{W}}_{i}^{Q},{\bm{E}}_{u}{\bm{W}}_{i}^{K},{\bm{E}}_{u}{\bm{W}}_{i}^{V})\end{split}$
where ${\bm{E}}_{u}$ is the input matrix computed from Equation 4, $h$ is a tunable hyper-parameter indicating the number of attention heads, ${\bm{W}}_{i}^{Q},{\bm{W}}_{i}^{K},{\bm{W}}_{i}^{V}\in\mathbb{R}^{d\times d/h}$ are the learnable weight matrices, and $\mathbf{W}^{O}\in\mathbb{R}^{d\times d}$ is also a learnable weight matrix. Attn is the attention function and is formally defined as:
(6) $\begin{split}\text{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})&=\text{softmax}(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d/h}})\mathbf{V}\end{split}$
The scale factor $\sqrt{d/h}$ is used to avoid large values of the inner product, which can lead to numerical instability.
Causality: In sequential recommendation, the prediction of the $L+1$ item should only depend on the first $L$ items that the user has interacted with in the past. However, the $L$-th output of the multi-head self-attention layer contains all the input information. Therefore, as in (Li et al., 2020; Kang and McAuley, 2018), we do not let the model attend to the future items by forbidding links between $Q_{i}$ and $K_{j}$ ($j>i$) in the attention function.
Point-Wise Feed-Forward Network: To add nonlinearity and interactions between different embedding dimensions, we follow previous works in sequential recommendation (Sun et al., 2019; Kang and McAuley, 2018; Li et al., 2020) and apply the same point-wise feed-forward network to the output of each multi-head self-attention layer. Formally, suppose the output of the multi-head self-attention layer is $\textbf{z}_{u}$, the point-wise feed-forward network is defined as:
(7) $\begin{split}\text{FFN}(\textbf{z}_{u})&=\text{ReLU}(\textbf{z}_{u}{\bm{W}}_{1}+\mathbf{b}_{1}){\bm{W}}_{2}+\mathbf{b}_{2}\end{split}$
where ${\bm{W}}_{1}\in\mathbb{R}^{d\times d}$ and ${\bm{W}}_{2}\in\mathbb{R}^{d\times d}$ are learnable weight matrices, and $\mathbf{b}_{1}\in\mathbb{R}^{d}$ and $\mathbf{b}_{2}\in\mathbb{R}^{d}$ are learnable bias vectors.
Stacking Layers: As shown in previous works (Kang and McAuley, 2018), stacking multiple multi-head self-attention layers and point-wise feed-forward networks can potentially lead to overfitting and instability during the training. Therefore, we follow previous works (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2020) and apply layer normalization (Ba et al., 2016) and residual connections to each multi-head self-attention layer and point-wise feed-forward network. Formally, we have:
(8) $g(\mathbf{x})=\mathbf{x}+\text{Dropout}(g(\text{LayerNorm}(\mathbf{x})))$
$g(\mathbf{x})$ is either the multi-head self-attention layer or the point-wise feed-forward network. Therefore, for every multi-head self-attention layer and point-wise feed-forward network, we first apply layer normalization to the input, then apply the multi-head self-attention layer or point-wise feed-forward network, and finally apply dropout and add the input $\mathbf{x}$ to the layer output. The LayerNorm function is defined as:
(9) $\text{LayerNorm}(\mathbf{x})=\alpha\odot\frac{\mathbf{x}-\mu}{\sqrt{\sigma^{2}+\epsilon}}+\mathbf{\beta}$
where $\odot$ denotes the element-wise product, $\mu$ and $\sigma$ are the mean and standard deviation of $\mathbf{x}$, $\alpha$ and $\mathbf{\beta}$ are learnable parameters, and $\epsilon$ is a small constant to avoid numerical instability.
4.1.6. Prediction
Given a sequence $\mathcal{S}_{u}$ of user $u$ as input, we denote $\mathbf{q}_{u}$ as the output of the popularity dynamics-aware transformer. Suppose at time $t^{+}$, we want to predict the likelihood of $v_{j}$ being the next item in the sequence, we first compute the item representation $\mathbf{e}_{j}^{t^{+}}$ from § 4.1.2. Then, we predict the score as the inner product of $\mathbf{q}_{u}$ and $\mathbf{e}_{j}^{t^{+}}$, formally:
(10) $\hat{y}(v^{t}_{j}|\mathcal{S}_{u})=<\mathbf{q}_{u},\mathbf{e}_{j}^{t^{+}}>$
Note that there is no information leakage in the prediction process, i.e., we do not assume access to the popularity statistics of $v_{j}$ at time $t^{+}$ if we are at time $t^{+}$ (§ 4.1.2).
4.2. Training Procedure
Now we present how to train the PrepRec model. Similar to (Kang and McAuley, 2018), we adopt the binary cross entropy loss as the objective function, formally:
(11) $\begin{split}\mathcal{L}=-\sum_{\mathcal{S}_{u}\in\mathcal{S}}\sum_{z\in[1,2,...,L-1]}&[\log\sigma(\hat{y}(v^{t}_{z+1}|\mathcal{S}_{u,:z}))\\ &+\sum_{j^{\prime}\notin\mathcal{S}_{u}}\log\sigma(1-\hat{y}(v^{t}_{j^{\prime}}|\mathcal{S}_{u,:z}))]\\ \end{split}$
where $\mathcal{S}_{u,:z}\coloneq\{v_{u,1},v_{u,2},...,v_{u,z}\}$. $v^{t}_{z+1}$ represents the $z+1$-th item in the sequence that happened at time $t$. We use Adam (Kingma and Ba, 2014) as the optimizer and train the model end-to-end. Note that compared to previous sequential recommenders, we do not have any parameters modeling item IDs. Essentially, we are only optimizing ${\bm{W}}_{p}$ and parameters related to the multi-head self-attention mechanism.
Dataset #users #items #actions avg length density
Office 101,133 27,500 0.74M 7.3 0.03%
Tool 240,464 73,153 1.96M 8.1 0.01%
Movie 70,404 40,210 11.55M 164.2 0.41%
Music 20,539 10,121 0.66M 32.2 0.32%
Epinions 30,989 20,382 0.54M 17.5 0.09%
Table 1. Dataset statistics
[…]
Metric R@10 N@10 R@10 N@10 R@10 N@10 R@10 N@10 R@10 N@10
Reference: Regular Sequential Recommendation
MostPop 0.450 0.272 0.459 0.274 0.586 0.361 0.519 0.327 0.438 0.296
BERT4Rec (Sun et al., 2019) $0.541^{*}$ $0.358^{*}$ 0.544 $0.350^{*}$ 0.900 0.728 $0.816^{*}$ $0.602^{*}$ 0.702 $0.512$
PrepRec 0.536 0.344 $0.551^{*}$ $0.339$ $0.908^{*}$ $0.738^{*}$ 0.782 0.573 $0.795^{*}$ $0.580^{*}$
[…]
Office $\rightarrow$ 0.540 0.326 0.838 0.624 0.755 0.542 0.724 0.512
Tool $\rightarrow$ 0.543 0.332 0.881 0.659 0.749 0.536 0.717 0.510
[…]
4.3. Zero-shot Inference
Suppose we are given a pre-trained model $\mathcal{F}$ trained on ${\bm{M}}$, where $\mathcal{F}$ is the scoring function learned from source domain ${\bm{M}}$. Denote the interaction matrix of the target domain as ${\bm{M}}^{\prime}$. We first compute the popularity dynamics of each item in ${\bm{M}}^{\prime}$ over a coarser period and a finer period. Then, we apply the pre-trained model $\mathcal{F}$ to ${\bm{M}}^{\prime}$ and compute the prediction score as:
(12) $\hat{y}(v^{t}_{j^{\prime}}|\mathcal{S}^{\prime}_{u})=\mathcal{F}(v^{t}_{j^{\prime}}|\mathcal{S}^{\prime}_{u},{\bm{M}}^{\prime})$
Note that in this procedure, we use the pre-trained model $\mathcal{F}$ trained on domain ${\bm{M}}^{\prime}$ to predict the next item $v^{t}_{j^{\prime}}$ that user $u^{\prime}$ will interact with in domain ${\bm{M}}^{\prime}$. We do not use any auxiliary information in either domain. In addition, none of the parameters in $\mathcal{F}$ are updated during the zero-shot inference process.
To summarize, in this section, we showed how to develop a pre-trained sequential recommender system based on the popularity dynamics of items. We enforce the structure of each interaction in the sequence by the positional encoding and introduce a relative time encoding for modeling time intervals between two consecutive interactions. In addition, we showed the training process and formally defined the zero-shot inference procedure. In the next section, we present experiments to evaluate PrepRec .
5. EXPERIMENTS
We present extensive experiments on five real-world datasets to evaluate the performance of PrepRec, following the problem settings in § 3. We introduce the following research questions (RQ) to guide our experiments: (RQ1) How well can PrepRec perform on zero-shot cross-domain and cross-application transfer? (RQ2) Why should we model popularity dynamics in sequential recommendation? (RQ3) What affects the performance of PrepRec ?
5.1. Datasets and Preprocessing
We evaluate our proposed method on five real-world datasets across different applications, with varying sizes, and density levels.
Amazon (Ni et al., 2019) is a series of product ratings datasets obtained from Amazon.com, split by product categories. We consider the Office and Tool product domains in our study. Douban (Song et al., 2019) consists of three datasets across different domains, collected from Douban.com, a Chinese review website. We work with the Movie and Music datasets. Epinions (Tang et al., 2012a, b) is a dataset crawled from product review site Epinions. We utilize the ratings dataset for our study.
We present dataset statistics in Table 1. We compute the density as the ratio of the number of interactions to the number of users times the number of items. Douban datasets (i.e., movie and music) are the densest and have no auxiliary information available, while the Amazon review datasets (i.e., office and tool) are the sparsest.
For fair evaluation, we follow the same preprocessing procedure as previous works (Kang and McAuley, 2018; Sun et al., 2019), i.e., we binarize the explicit ratings to implicit feedback. In addition, for each user, we sort interactions by their timestamp and use the second most recent action for validation, the most recent action for testing, and the rest for training.
5.2. Baselines and Experimental Setup
Baselines: Our baselines (supplementary materials contain detailed descriptions) include classic general recommendation models (e.g., MostPop, BPR (Rendle et al., 2009), NCF (He et al., 2017), LightGCN (He et al., 2020)) and state-of-the-art sequential recommendation models (e.g., Caser (Tang and Wang, 2018), SasRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019), TiSasRec (Li et al., 2020), CL4SRec (Xie et al., 2022)).
Dataset Office Tool Movie Music Epinions
Metric R@10 N@10 R@10 N@10 R@10 N@10 R@10 N@10 R@10 N@10
General Recommender Systems
MostPop 0.450 0.272 0.459 0.274 0.586 0.361 0.519 0.327 0.438 0.296
BPR (Rendle et al., 2009) 0.457 0.289 0.363 0.216 0.747 0.477 0.646 0.434 0.568 0.397
NCF (He et al., 2017) 0.446 0.266 0.388 0.239 0.784 0.505 0.652 0.437 0.570 0.396
LightGCN (He et al., 2020) 0.465 0.293 0.463 0.275 0.793 0.512 0.665 0.447 0.575 0.396
Sequential Recommender Systems
Caser (Tang and Wang, 2018) 0.512 0.334 0.496 0.297 0.891 0.701 0.796 0.576 0.674 0.475
SasRec (Kang and McAuley, 2018) 0.539 0.354 0.536 0.337 0.918 0.749 0.816 0.599 $0.705^{*}$ 0.501
BERT4Rec (Sun et al., 2019) 0.541 0.358 0.544 0.350 0.900 0.728 $0.816^{*}$ $0.602^{*}$ 0.702 $0.512^{*}$
TiSasRec (Li et al., 2020) 0.531 0.349 0.539 0.341 $0.918^{*}$ $0.752^{*}$ 0.809 0.523 0.701 0.499
CL4SRec (Xie et al., 2022) $0.550^{*}$ $0.358^{*}$ $0.548^{*}$ $0.352^{*}$ 0.899 0.725 0.813 0.597 0.662 0.481
PrepRec 0.536 0.344 0.551 0.339 0.908 0.738 0.782 0.573 0.795 0.580
PrepRec $\Delta$2.5% -3.9% +1.2% -3.6% -1.1% -1.8% -1.9% -4.8% +12.7% +13.3%
Interp 0.648 0.483 0.659 0.482 0.929 0.769 0.851 0.653 0.816 0.640
Interp $\Delta$ +17.8% +34.9% +20.3% +35.0% +1.1% +2.3% +4.3% +8.5% +15.7% +25.0%
Table 3. Regular sequential recommendation results, RQ2, (§ 5.4.1). We make bold the best results and mark the best baseline results with ${}^{\prime}*^{\prime}$. Interp represents the interpolation results between PrepRec and BERT4Rec. PrepRec $\Delta$ denotes the performance difference between PrepRec and the best results among the selected baselines, similar for Interp $\Delta$. PrepRec achieves comparable performance to the state-of-the-art sequential recommenders, with only on average 0.2% worse than the best performing sequential recommenders in R@10 while having only a fraction of the model size (Table 5). After a simple post-hoc interpolation, we outperform the state-of-the-art sequential recommenders by 11.8% in R@10 on average.
Following previous works (Koren, 2008; He et al., 2017; Kang and McAuley, 2018; Sun et al., 2019), we adopt the leave-one-out evaluation method: for each user, we pair the test item with 100 unobserved items according to the user’s interaction history. Then we rank the test item for the user among the 101 total items. We use two standard evaluation metrics for top-$k$ recommendation: Recall@k (R@k) and Normalized Discounted Cumulative Gain@k (N@k). Our model explicitly utilizes popularity information. Therefore, we also present results where we sample the negatives based on their popularities, i.e., popular items have higher probabilities of being sampled as negatives. We report the average of R@k and N@k over all the test interactions.
We use publicly available implementations for the baselines. For fair evaluation, we set dimension size $d$ to 50, max sequence length $L$ to $200$, and batch size to $128$ in all models. We use an Adam optimizer and tune the learning rate in the range $\{10^{-4},10^{-3},10^{-2}\}$ and set the weight decay to $10^{-5}$. We use the dropout regularization rate of $0.3$ for all models. We set $\gamma=0.5$ in Equation 1, whose reason we will discuss the reason in supplement materials. We define the coarse and fine period to be 10 and 2 days respectively, and we fix the window size to be $m=12$ and $n=4$ for all datasets (§ 4.1.2). We train PrepRec for a maximum of 80 epochs. All experiments are conducted on a Tesla V100 using PyTorch. We repeat each experiment 5 times with different random seeds and report the average performance.
Figure 3. Zero-shot Transfer Results (R@10) with Gaussian noise added to the item popularity statistics (§ 5.3.2). We find that PrepRec is relatively robust to noise.
5.3. Zero-shot Transferability (RQ1)
5.3.1. Zero-shot Transfer Results
We follow the zero-shot inference setting introduced in § 4.3 and report the results in Table 2. We also include the results of PrepRec and the best-performing sequential recommenders trained on the target dataset for reference. In the zero-shot setting, PrepRec shows minimal performance reduction in the target datasets (i.e., 6% maximum and 2% average reduction in R@10). The best zero-shot transfer results from PrepRec only fall short against the selected sequential recommendation baselines by up to 4% and even outperform (by up to 6.5%) them on the Epinions and Office. We found that PrepRec trained on douban-movie and amazon-tools show the highest generalizability, even outperforming the target-trained models on Music (0.811 vs. 0.782 on R@10). We conjecture that this is because Movie is the largest dataset in terms of the number of interactions. Overall, these results show PrepRec ’s effectiveness in zero-shot transfer without any training on interaction data or side information. In addition, this experiment also demonstrates that the popularity dynamics-based item and sequence representations are generalizable across domains.
5.3.2. Robustness to Noise
We further investigate the robustness of PrepRec to possible noise in zero-shot transfer by adding Gaussian noise $\sim\mathcal{N}(0,\sigma)$ to the item popularity statistics and evaluate the zero-shot transfer performance on Douban-Music and Epinions from Douban-Movie. We randomly choose some percentage of items in the sequence to add noise, as indicated in  Figure 3. We find that PrepRec is relatively robust to noise, maintaining robust performance across different noise levels at 20% noised interaction. We attribute this to the model’s ability to learn from the overall popularity dynamics, which is less affected by noise in individual item popularity statistics. In addition, when the noise level is relatively low, e.g., $\sigma\leq 5$, even if 100% of the sequence is noised, PrepRec still holds the performance, indicating significant item popularity shifts exist in the sequence (Figure 1).
Dataset Music Office Epinions
Metric R@10 N@10 R@10 N@10 R@10 N@10
MostPop 0.197 0.139 0.099 0.046 0.163 0.110
SasRec (Kang and McAuley, 2018) 0.749 0.519 0.453 0.291 0.658 0.442
BERT4Rec (Sun et al., 2019) 0.747 0.519 0.461 0.299 0.655 0.456
PrepRec 0.739 0.523 0.443 0.280 0.762 0.551
PrepRec $\Delta$1.3% +0.7% -2.2% -6.3% +15.8% +17.2%
Table 4. Regular sequential recommendation results (§ 5.4.1) with popularity-based negative sampling. PrepRec can learn discriminative item and sequence representations despite depending only on popularity statistics.
5.4. Why Popularity Dynamics? (RQ2)
PrepRec shows excellent performance in zero-shot sequential recommendation. Therefore, we ask: what is the role of popularity dynamics in sequential recommendation, and how much does it explain the performance of state-of-the-art sequential recommenders? Therefore, we propose the following experiments to investigate the importance of popularity dynamics in sequential recommendation.
5.4.1. Regular Sequential Recommendation (RQ2)
We show comparisons of PrepRec against state-of-the-art sequential recommenders in the regular sequential recommendation tasks (Table 3), i.e., all models are trained from scratch. PrepRec achieves competitive performance—within 2% in R@10 and 5% in N@10, with the state-of-the-art baselines. On Epinions, PrepRec even outperforms all baselines by 7.3%, particularly impressive since PrepRec has significantly fewer model parameters (Table 5).
Dataset Office Tool Movie Music Epinions
SasRec 1.331M 3.581M 2.044M 0.542M 1.054M
BERT4Rec 2.687M 7.233M 4.126M 1.094M 2.127M
TiSasRec 1.367M 3.617M 2.127M 0.578M 1.09M
PrepRec 0.045M 0.045M 0.045M 0.045M 0.045M
Table 5. Comparison of model sizes (i.e., number of learnable parameters in millions) over different datasets. PrepRec is 12 to 90x smaller.
PrepRec explicitly models popularity information and the MostPop demonstrates decent performance compared to the remaining baselines, thus we conduct an additional experiment (Table 4) where we sample the unobserved (negative) items based on their popularity (Sun et al., 2019). As shown in Table 4, MostPop’s performance dropped significantly, while PrepRec shows even more competitive performance on some datasets (e.g., Music and Epinions). This suggests PrepRec learns discriminative item and sequence representations.
PrepRec learns item representations through popularity dynamics, which is conceptually different from learning representations specific to each item ID. Therefore, we propose a simple post-hoc interpolation to investigate how much can popularity dynamics explain the performance of state-of-the-art sequential recommenders. We interpolate the scores from PrepRec with the scores from BERT4Rec as follows: $\hat{y}_{intp}(v^{t}_{j}|\mathcal{S}_{u})=\alpha*\hat{y}_{O}(v^{t}_{j}|\mathcal{S}_{u})+(1-\alpha)*\hat{y}_{S}(v_{j}|\mathcal{S}_{u})$, where $\hat{y}_{O}(v^{t}_{j}|\mathcal{S}_{u})$ and $\hat{y}_{S}(v_{j}|\mathcal{S}_{u})$ are the scores from PrepRec (Equation 10) and BERT4Rec, respectively. We set $\alpha=0.5$ for all datasets. After interpolation, the performance significantly boosts by up to 34.9% in N@10. Gains are largest in the medium and low-density datasets (Epinions, Amazon), indicating that our model complements existing methods in sparse datasets where item embeddings are less informative. Therefore, it is crucial to consider popularity dynamics to maximize performance.
5.4.2. Qualitative Analysis on Regular Sequential Recommendation
Figure 4. Recommendation results for different item popularity groups (§ 5.4.2), where Group 1 represents the least popular items, and Group 5 represents the most popular items. PrepRec achieves better performance on long-tail items while having competitive performance on popular items.
We analyze the performance of PrepRec in detail. We separate test items into equally sized groups based on their popularity in the training set then compute the average R@10 and N@10 for each group (Figure 4). PrepRec achieves better performance on item group with the least interactions, i.e., long-tail items, while the SasRec and BERT4Rec show stronger performance on popular items. Long-tail item recommendation is a particularly challenging task explored by many previous works (Sankar et al., 2021) and requires recommenders able to learn high-quality representations with just a few interactions. This corresponds to our observation that PrepRec is more robust to data sparsity and can learn discriminative item and sequence representations (§ 5.4.1), showing that long-tail item recommendation can benefit from PrepRec ’s popularity dynamics-based item representations.
5.5. What affects PrepRec performance? (RQ3)
5.5.1. Ablation Study
Dataset Music Office Epinions
Metric R@10 N@10 R@10 N@10 R@10 N@10
PrepRec 0.782 0.573 0.536 0.344 0.795 0.580
w/o Relative Time ${\bm{T}}$ (§ 4.1.3) 0.734 0.514 0.541 0.334 0.782 0.562
w/o Positional ${\bm{P}}$ (§ 4.1.4) 0.765 0.544 0.530 0.332 0.772 0.554
w/o Popularity Dynamics $\mathcal{P}$ 0.800 0.594 0.530 0.341 0.761 0.560
w/o Popularity Dynamics $\mathcal{H}$ 0.705 0.582 0.525 0.337 0.730 0.533
Sinuoisal Popularity Encoding 0.779 0.570 0.529 0.340 0.772 0.561
Table 6. Ablation study of PrepRec ’s different variants.
Here, we assess the importance of different components crucial to PrepRec , i.e., relative time encoding (§ 4.1.3), positional encoding (§ 4.1.4), popularity encoder $E_{p}$ (§ 4.1.1), and resolutions for popularity dynamics (§ 4.1.2). We find that removing relative time encoding ${\bm{T}}$ results in the largest performance drop on both the Music and Office datasets. This suggests that the relative time encoding is crucial for effectively capturing the popularity dynamics. Removing positional encoding ${\bm{P}}$ results in a maximum of 2.2% drop in R@10 on the Office dataset, indicating positional encoding is important for capturing sequential information. In addition, changing $E_{p}$ to the non-linear sunusoidal encoding shows worse performance on all datasets, meaning that the linear encoding is more suitable for capturing the popularity dynamics. On the music dataset, removing coarse popularity encoding $\mathcal{P}$ results improves the performance by 2% in R@10, while removing fine popularity encoding $\mathcal{H}$ results in a 7.5% drop in R@10. This suggests that the music domain is more sensitive to recent trends in popularity. Coarse and fine popularity encodings complement each other on other datasets.
Dataset Music Office Epinions
Metric R@10 N@10 R@10 N@10 R@10 N@10
$\gamma=0$ (Curr-pop) 0.749 0.542 0.512 0.328 0.689 0.496
$\gamma=0.25$ 0.764 0.529 0.538 0.338 0.761 0.562
$\gamma=0.5^{*}$ (weighted -pop) 0.782 0.573 0.536 0.344 0.795 0.580
$\gamma=0.75$ 0.755 0.520 0.543 0.336 0.747 0.519
$\gamma=1$ (cumul-pop) 0.695 0.452 0.530 0.330 0.733 0.505
Table 7. Recommendation results for varying the discounting factor $\gamma$ in § 4.1.2. $\gamma=0.5$ is the default setting, denoted by ${}^{\prime}*^{\prime}$. We find that $\gamma=0.5$ generally outperforms the other two settings
5.5.2. Effect of discounting factor $\gamma$
We examine the effect of different preprocessing weights $\gamma$ used in popularity calculation (§ 4.1.1). In particular, $\gamma=1$ corresponds to the cumulative popularity, or in other words, at a given time period $t$, the overall number of interactions up to period $t$. On the other hand, $\gamma=0$ corresponds to the current popularity, or percentiles are calculated over interactions just in $t$, same as $b_{j}^{t}$ in Equation 1. When $\gamma=0.5$, it can be interpreted as interactions being exponentially weighted by time, with a half-life of 1 time period. We find that $\gamma=0.5$ outperforms the other two settings, with the largest gains of around 12% R@10 and 27% N@10 over cumul-pop in the dense Music dataset, and the largest gains over curr-pop in the sparser Office (5% N@10 and 4% N@10) and Epinions ($15\%$ R@10 and 17%N@10) datasets. We suspect this is due to cumulative measures in denser datasets failing to capture recent trends due to the large historical presence, while current-only measures in sparser datasets convey too little or noisy information and lose the information of long-term trends. curr-pop shows decent performance on the Music dataset, suggesting that Music trends might be more cyclical and thus the current popularity is more informative.
5.5.3. Effect of Different Time Horizons
Dataset Music Office Epinions
Metric R@10 N@10 R@10 N@10 R@10 N@10
Fine:2 days; Coarse:10 days 0.782 0.573 0.536 0.344 0.795 0.580
Fine:4 days; Coarse:15 days 0.778 0.553 0.537 0.341 0.790 0.574
Fine:7 days; Coarse:30 days 0.760 0.509 0.526 0.334 0.757 0.543
Table 8. Recommendation results for varying time horizons. Fine and coarse time horizons are used for short-term and long-term popularity dynamics respectively (§ 4.1.1).
We study the effect of different time horizons to PrepRec . We found that in general, long-term horizons of 30 days and short-term horizons of 7 days perform worse than the other settings. This is likely because the long-term horizon might lead to the lack of resolutions in popularity statistics. We also find that depending on the dataset, the effect of different time horizons also varies. For example, both Music and Epinions show larger performance decrease from short to long-term horizons than Office. This could be because Music and Epinions are more sensitive to recent trends than Office, or their data are denser in terms of time granularity.
5.6. Fine-tune Capability
[…]
Metric R@10 N@10 R@10 N@10 R@10 N@10
PrepRec 0.803 0.591 0.472 0.300 0.489 0.264
SasRec 0.815 0.599 0.437 0.290 0.433 0.245
BERT4Rec 0.816 0.602 0.407 0.249 0.433 0.255
Table 9. Recommendation results for fine-tuning PrepRec. We fine-tune PrepRec and retrain the baselines from scratch on the target dataset.
We also investigate PrepRec’s fine-tune capability. To ensure target datasets are smaller than the source, we further process the target datasets such that they are no more than 10% of the source datasets’ total interactions. After further processing, we follow the same experimental setup in § 5.2. We fine-tune PrepRec and retrain the baselines from scratch on the target dataset and report the results in Table 9. We find that PrepRec , after fine-tuning, outperforms the selected baselines on Office and Epinions by up to 12.9%, indicating that PrepRec is capable of learning from the limited data and can be further fine-tuned to achieve better performance.
5.7. Discussion
PrepRec demonstrates the strong ability for zero-shot transfer. We argue that PrepRec is particularly useful in the following scenarios: 1) initial sequential model when the data in the domain is sparse; 2) backbone for developing more complex sequential recommenders (i.e., prediction interpolation) 3) online recommendation settings.
PrepRec captures the popularity shifts in the sequence and is complementary to state-of-the-art sequential recommenders. It is worth noting that item popularity dynamics might not capture everything in users’ preferences, but we believe they are orthogonal components towards capturing user preferences, which could explain why the interpolation results substantially outperform both PrepRec and the selected state-of-the-art baselines (Table 3).
Additionally, time granularity is also crucial for popularity dynamics, and sequence analysis requires careful consideration of the time horizon. Intuitively, when the dataset time precision is less accurate, i.e., weeks or days, we expect the performance to decrease as the sequential information and popularity dynamics become muddled. If the time precision in the training data increases, we can expect more accurate user sequences and more accurate measures of popularity dynamics. In general, time precision will not significantly impact the performance of PrepRec in most scenarios as in practice, online platforms can record precise time data for each user-item interaction. We will include more discussion in the arXiv version of this paper (Wang et al., 2024).
6. CONCLUSION
In this paper, using the critical insight of popularity dynamics in the user’s sequence, we developed a novel pre-trained sequential recommendation framework, PrepRec, for the zero-shot, cross-domain setting without any auxiliary information. PrepRec learned transferable, universal item representations via popularity dynamics-aware transformers. We empirically showed that PrepRec can achieve excellent zero-shot transfer to a target domain, comparable to state-of-the-art sequential recommenders trained on the target domain. With extensive within-domain experiments, we found performance gains of 11.8% when we interpolated PrepRec ’s results with state-of-the-art sequential recommenders, indicating that PrepRec is learning complementary information. We posit that popularity dynamics are crucial for developing generalizable sequential recommenders.
As part of future work, we plan to investigate: 1) developing more complex sequential recommenders by using PrepRec as a backbone (i.e., prediction interpolation and auxiliary information), and 2) exploring online recommendation settings.
Acknowledgements.
This work was generously supported by the National Science Foundation (NSF) under grant number 2312561. We also would like to thank the anonymous reviewers for their valuable feedback.